### v1

# Open-World Skill Discovery from Unsegmented Demonstrations

Abstract
Learning skills in open-world environments is essential for developing agents capable of handling a variety of tasks by combining basic skills. Online demonstration videos are typically long but unsegmented, making them difficult to segment and label with skill identifiers. Unlike existing methods that rely on sequence sampling or human labeling, we have developed a self-supervised learning-based approach to segment these long videos into a series of semantic-aware and skill-consistent segments. Drawing inspiration from human cognitive event segmentation theory, we introduce Skill Boundary Detection (SBD), an annotation-free temporal video segmentation algorithm. SBD detects skill boundaries in a video by leveraging prediction errors from a pretrained unconditional action-prediction model. This approach is based on the assumption that a significant increase in prediction error indicates a shift in the skill being executed. We evaluated our method in Minecraft, a rich open-world simulator with extensive gameplay videos available online. Our SBD-generated segments improved the average performance of conditioned policies by 63.7% and 52.1% on short-term atomic skill tasks, and their corresponding hierarchical agents by 11.3% and 20.8% on long-horizon tasks. Our method can leverage the diverse YouTube videos to train instruction-following agents. The project page can be found in https://craftjarvis.github.io/SkillDiscovery/.
[…]
Yitao Liang <yitaol@pku.edu.cn>
[…]
Jingwen Deng <dengjingwen@stu.pku.edu.cn>, Zihao Wang <zhwang@stu.pku.edu.cn>, Anji Liu <liuanji@cs.ucla.edu>
1 Introduction
Method Illustration Type
Sequential sampling [Uncaptioned image] rule-based
Reward-driven [Uncaptioned image] rule-based
Top-down [Uncaptioned image] rule-based
Bottom-up [Uncaptioned image] rule-based
Ours (SBD) [Uncaptioned image] learning-based
Table 1: Comparisons between existing segmentation methods and our method SBD. Existing methods usually rely on human-designed rules, while our method is learning-based. Sequential sampling can result in a single skill spanning different segments or multiple skills located within one segment. Reward-driven methods require additional reward information, which is challenging for human annotators to label. Top-down methods often result in limited skill diversity and high computation costs. Bottom-up methods are limited in fully observable environments and hard in visually partially observable environments. [Uncaptioned image] are visual observations and [Uncaptioned image] are actions.
Most existing LLM-based instruction-following agents adopt a two-layer structure of planner and controller to complete tasks, which first convert the instruction into atom skills through the planner, and then use a conditioned policy to convert it into actions based on the current observation to interact with the environment (Wang et al., 2025, 2023b, 2024b; Li et al., 2024; Wang et al., 2024a; Cheng et al., 2024). When training these agents, we need to first convert long sequences into a series of short atom skill sequences, and then train the controller and planner separately. However, interactions in real-world videos are often long-term and unsegmented with detailed language skills.
Learning skills from long-sequence videos is critical for building such hierarchical agents. However, the concept of a “skill” is ill-defined and varies widely across domains such as video gaming (Cai et al., 2023b; Wang et al., 2024a; Hafner et al., 2023; Zhang et al., 2024), robotic control (Zitkovich et al., 2023), and autonomous driving (Wang et al., 2023a). This ambiguity, along with the immense diversity of skills in open worlds, makes skill learning particularly challenging, especially in partially observable settings. To enable open-world skill learning from unsegmented demonstrations (continuous streams of observation-action pairs without explicit labels), the first challenge is to segment these streams into semantically meaningful, self-contained skills.
We list the existing segmentation methods in Table 1, including the plain sequential sampling, reward-driven methods, top-down, and bottom-up methods. The native sequential sampling segmentation methods (Cai et al., 2023b; Lifshitz et al., 2023), divide videos into segments of predefined lengths (e.g., fixed length, uniform distribution). However, these methods do not ensure that each segment contains a distinct skill. Additionally, predefined lengths may not match the actual distribution of skill lengths in real-world scenarios (see Section 4.4). Reward-driven methods (Sutton et al., 1999) discover skills through the environment’s reward signal. It is limited by its inability to capture skills that lack associated rewards and by the risk of splitting a single skill into multiple segments when rewards are repeatedly gained during execution. Top-down methods (Shiarlis et al., 2018; Belkhale et al., 2024) rely on predefined skill sets from human experts. They use manual labeling or supervised learning to segment videos. Although this approach can produce reasonable results, it is expensive and limited by the narrow range of predefined skills. Bottom-up methods (Zhu et al., 2022; Pertsch et al., 2025) use algorithms such as agglomerative clustering or byte-pair encoding (BPE) (Gage, 1994) to split action sequences. However, they struggle in partially observable settings where both observations and actions must be considered because these algorithms cannot deal with observations due to the high dimensionality. All the above methods usually rely on human-designed rules to segment the unsegmented videos, which highlights the need for an effective and label-free method that can automatically segment skills from unsegmented demonstrations in open worlds.
Inspired by event segmentation theories (EST) (Zacks et al., 2007), which propose that humans naturally partition continuous experiences into discrete events when prediction errors in perceptual expectations rise, we introduce Skill Boundary Detection (SBD)—a method for autonomously identifying potential skill boundaries within extended trajectories. SBD employs a predictive model trained on a dataset of unsegmented videos to forecast future actions based on past observations, effectively capturing temporal dependencies (Baker et al., 2022). We then use this pretrained model to make predictions on an unsegmented video and compare them to the ground truth actions. A significant increase in prediction error indicates a shift in the skill being executed (Theorem 3.4), enabling us to detect boundaries between different skills in the video. SBD relies on self-supervised learning, removing the need for additional human labeling. This allows SBD to utilize a wide range of YouTube videos to train instruction-following agents.
We evaluate SBD in open-world Minecraft (Guss et al., 2019; Fan et al., 2022; Lin et al., 2023). First, we apply SBD to create a segmented Minecraft video dataset. We then train the video-conditioned policy (Cai et al., 2023b) and the language-conditioned policy (Lifshitz et al., 2023) on this dataset. We evaluate these policies on diverse Minecraft skills benchmark (Lin et al., 2023). The results show significant improvements compared to the existing policies (Cai et al., 2023b; Lifshitz et al., 2023), with performance increases of 63.7% and 52.1% compared to the original methods, respectively. We also test their corresponding hierarchical agents using behavior cloning and in-context learning, achieving performance increases of 11.3% and 20.8%, respectively. These findings underscore the effectiveness of our proposed method for skill discovery from unsegmented demonstrations and its potential to advance open-world skill learning.
Figure 1: Pipeline of our method SBD for discovering skills from unsegmented demonstration videos. Stage I: An unconditional Transformer-XL based policy model (Dai et al., 2019; Baker et al., 2022) is pretrained on an unsegmented dataset to predict future actions (labeled by an inverse dynamics model) based on past observations. Stage II: The pretrained unconditional policy will produce a high predicted action loss when encountering uncertain observations (e.g., deciding whether to kill a new sheep) in open worlds. These timesteps should be marked as skill boundaries, indicating the need for additional instructions to control behaviors. We segment the long unsegmented videos into a series of short atomic skill demonstrations. Stage III: We train a conditional Transformer-XL based policy model on the segmented dataset to master a variety of atomic skills. Stage IV: Finally, we use hierarchical methods (a combination of vision-language models and the conditional policy) to model the long demonstration and follow long-horizon instructions.
2 Problem Formulation
Skills in behavioral cloning
Behavioral cloning learns a policy $\pi$ that maps observations $o$ from the observation space $\mathcal{O}$ to actions $a$ in the action space $\mathcal{A}$ using supervised learning on observation-action pairs $(o_{i},a_{i})_{i=1}^{T}$ obtained from expert demonstrations. In partially observable environments, past observations $o_{1:t}$ are often used instead of the current observation $o_{t}$ as input to the policy $\pi$. Additionally, to allow the agent to dynamically adjust its behavior based on its objective, recent advances (Khandelwal et al., 2021; Du et al., 2021; Xie et al., 2023) focus on learning goal-conditioned policies $\pi(o_{1:t},g)$, where $g$ represents the goal. This type of goal-conditioned policy can be considered a “skill”.
Hierarchical agent architecture
Directly learning policy $\pi(o_{1:t},g)$ is challenging in complex open-world environments like Minecraft. Goals such as build a house or mine diamond blocks involve numerous intermediate steps, making it computationally and conceptually difficult to model an agent’s behavior without breaking the goal into manageable parts. To tackle this, hierarchical agent architectures are often used (Wang et al., 2024a; Driess et al., 2023; Cai et al., 2024a; Iqbal et al., 2022). These architectures consist of a high-level planner and a low-level controller. The planner decomposes the overall goal into a series of sub-goals and decides which one the controller should address. The controller then focuses on achieving the current sub-goal by employing a specific skill. Specifically, for a certain $t$, the policy can be represented as:
$\pi(a|o_{1:t},g)=\pi_{1}(g_{t}|o_{1:t},g)\pi_{2}({a|o_{1:t},g_{t}})$ (1)
where $\pi_{1}$ is the planner’s policy and $\pi_{2}$ is the controller’s policy. The planner periodically (not at every step) checks whether the current sub-goal has been finished and determines whether to change policy, so $g_{t}$ is relatively static and does not change frequently.
Identifying implicit skills in offline trajectories
A critical challenge in training hierarchical agents, as described above, is the identification of skills within the offline trajectory data. To be more specific, given $\tau=(o_{i},a_{i})_{i=1}^{T}$, we want to split it into $\tau_{1}=(o_{i},a_{i})_{i=1}^{t_{1}},\tau_{2}=(o_{i},a_{i})_{i=t_{1}}^{t_{2}},\ldots,\tau_{k}=(o_{i},a_{i})_{i=t_{k-1}}^{T}$ where each sub-trajectories $\tau_{j}$ shows an independent skill $\pi(g_{j})$. Since some controllers use self-supervised methods to learn sub-goals (Lifshitz et al., 2023; Cai et al., 2023b), we can ignore the language labels of goal $g_{j}$, but must identify $t_{j}$ (i.e., the moments at which the agent adopts a different skill). For clarity, in the subsequent sections, we denote the skill adopted at a certain time step $t$ as $\pi_{t}$, corresponding to $\pi(g_{t})$.
3 Method
We first introduce our overall pipeline to build policies and long-horizon hierarchical agents based on unsegmented videos in Section 3.1. Then we introduce how we split the unsegmented demonstration into our policy and agent learning sequences in Section 3.2. We finally introduce the implementation details on Section 3.3.
3.1 Pipeline
Given a long sequence of trajectories $D={(o_{i},a_{i})}_{i=1}^{T}$ with interleaved visual observations $o_{i}$ and actions $a_{i}$, we first train an unconditioned policy $\pi_{\text{uncondition}}$ using imitation learning to predict the actions as follows:
$\min_{\theta}\sum_{t\in[1\dots T]}-\log\pi_{\text{uncondition}}(a_{t}|o_{1},\dots,o_{t}).$ (2)
When training on action label-free YouTube videos, the action labels are generated from Inverse Dynamics Model (Baker et al., 2022). We then evaluate this unconditioned policy $\pi_{\text{uncondtion}}$ on the entire dataset $D$ of long sequences. We developed an automated method called SBD to segment the complete video into a series of short segments $D_{seg}={(o_{i},a_{i})}_{i=m}^{n}$ based on these evaluation results, where $m$ and $n$ is the selected segmentation timestamps. These short segments $D_{seg}$ are then used for skill learning to obtain conditioned policies (Cai et al., 2023b; Lifshitz et al., 2023). Finally, we integrate these skills with vision language models to build hierarchical agents, enabling them to learn and follow long-horizon tasks in long video sequences (Wang et al., 2025, 2024b). We show the overall pipeline is illuminated in Figure 1.
3.2 Skill Boundary Detection
Algorithm 1 Skill Boundary Detection
1:  Input: $(o_{i},a_{i},e_{i})_{i=1}^{T}$, a model $M$ to predict $a_{t}$ given $o_{1:t}$. $e_{i}$ is the boolean external information indicating whether this step should be a boundary.
[…]
3:  for $t\leftarrow 1$ to $T$ do
4:     $\text{loss}\leftarrow M(a_{t}\mid o_{\text{begin}:t})$
[…]
8:        $\text{begin}\leftarrow t$
[…]
10:     end if
11:  end for
12:  Return: boundaries
SBD takes a long unsegmented trajectory as input and learns an unconditional model to predict the next action based on previous observations. The model does not know about implicit skill transitions. To simulate the model’s memory, a sliding window is used. At each time step $t$, the model predicts the next action and compares it with the ground truth to compute the loss. A skill transition is considered likely if the loss exceeds the average loss by a hyperparameter, GAP, or if an external indicator is true. In such cases, $t$ is marked as a boundary. The model’s memory is then cleared, and the algorithm proceeds to analyze the next sub-trajectory. The two core components of the algorithm are loss and external information.
3.2.1 Boundary with Entropy Loss
In this section, we explain the core idea of our method: why loss can indicate skill boundaries. To support this, we introduce three key assumptions about skills: skill consistency, skill confidence, and action deviance at skill transition.
The first assumption is that the skill being used does not change frequently. The parameter $K$ is a large number that determines how rare these changes are. The idea is that the agent should consistently stick to a skill unless there is a strong reason to switch to another, which would result in an increase in predictive loss for detecting a skill transition.
Assumption 3.1 (Skill Consistency).
There exists an adequately large $K$, $\forall t,$
$P(\pi_{t+1}\neq\pi_{t}|o_{1:t+1})<1/K$ (3)
The second assumption states that, at any given time, the agent is confident in the action it takes. This means the probability of the agent choosing an action with high confidence is very high. The parameter $c$ sets the minimum confidence level, while $\delta$ is a small value representing the frequency with which the policy might act with lower confidence. This assumption ensures that the policy reliably makes decisions based on its learned skills, rather than being uncertain in its actions. This helps the unconditional model make accurate predictions when the agent does not change its skill.
Assumption 3.2 (Skill Confidence).
There exists $c$ and an adequately small $\delta,\forall t,$
$P(\pi_{t}(a_{t}|o_{1:t})>c)>1-\delta$ (4)
The third assumption posits that when an agent changes its skill, it is likely to perform an action that is significantly less probable under the previous skill. This "surprising" action generates a clear signal, enabling us to detect the skill transition. Although the agent might change policies without performing a surprising action, such instances suggest that the agent is in a transitional, ambiguous phase between two policies, which is inherently challenging to identify. In essence, we aim to recognize the first clear skill boundary.
Assumption 3.3 (Action Deviance at Skill Transition).
There exists $m$, $\forall t$, when $\pi_{1}=\pi_{2}=...=\pi_{t}\neq\pi_{t+1}$,
$\frac{\pi_{t}(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}\pi_{t}(a_{i}|o_{1:i}))^{1/t}}<\frac{1}{2}m$ (5)
By the law of total probability, we have
$\begin{split}&P(a_{t+1}|o_{1:t+1})=P(\pi_{t+1}=\pi_{t}|o_{1:t+1})\pi_{t}(a_{t+1}|o_{1:t+1})\\ &\quad+\sum_{\pi\neq\pi_{t}}P(\pi_{t+1}=\pi|o_{1:t+1})\pi(a_{t+1}|o_{1:t+1})\end{split}$ (6)
Intuitively, if the agent changes its skill, the predictive probability will be low. This is because, in LABEL:eq:transition, in the first term the probability of $a_{t+1}$ under the original skill $\pi_{t}(a_{t+1}|o_{1:t+1})$ is low, and in the second term the probability of skill transition is low. On the contrary, if the agent does not change its skill, both terms will be high so the predictive probability will be high. Therefore, we have the following theorem regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
Theorem 3.4 (bounds of relative predictive probability).
If $\pi_{1}=\pi_{2}=...=\pi_{t}=\pi_{t+1}$, then we have
$P(\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}>\frac{(K-1)c}{K})>1-\delta$ (7)
If $\pi_{1}=\pi_{2}=...=\pi_{t}\neq\pi_{t+1}$, then we have
$\begin{split}P(\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}<\frac{Km}{2(K-1)}+\frac{1}{c(K-1)})\\ >1-t\delta\end{split}$ (8)
Proof.
See Appendix A for the detailed proof. ∎
If $c>m$ and $(K-4)c^{2}>2$, then the lower bound of relative predictive probability under skill transition in Theorem 3.4 is greater than the upper bound of it under skill non-transition. Therefore, the theorem demonstrates that the model predicts actions less accurately when the agent changes its skill. Our action-prediction model uses the negative log-likelihood of actions $-\log P(a_{t}|o_{1:t})$ as loss. Consequently, at line 6 of Algorithm 1, when the loss exceeds the average loss by a predefined threshold, it is likely that a skill transition has occurred.
3.2.2 Boundary with Auxiliary Information
In Section 3.2.1, we demonstrate that our algorithm can identify the first distinguishable skill boundary using only observations and actions. However, some datasets include additional external information, such as privileged in-game data or VLM detections. Incorporating these auxiliary signals improves the detection of skill boundaries that are otherwise difficult to recognize through loss-based boundary detection alone. For example, in Minecraft, crafting an item does not involve abrupt changes in mouse movement, making it hard to detect boundaries purely from predictive action loss. In such cases, in-game logs of successful crafting events provide valuable hints. It is important to note that this is an optional component of our algorithm; as we will discuss in Section 4.3, it also performs well on datasets without external auxiliary information.
3.3 Implemetation Details
Unconditioned Policy and Segmentation Details
For the unconditioned policy $\pi_{\text{uncondition}}$ in Stage I, we use the pre-trained vpt-3x models (Baker et al., 2022). This is a foundational Transformer-XL (Dai et al., 2019) based model pretrained on large-scale YouTube data and finetuned on the early-game dataset using behavioral cloning. The hyperparameter GAP in Stage II is set to 18 considering the average trajectory length and semantics. As for the external information, we use event-based information such as mine_block:oak_log or use_item:torch available in the contractor dataset (Baker et al., 2022). To identify potential skill transitions, only the final event in a series of identical consecutive events is marked as positive, as it signifies a possible skill transition. For instance, if the agent chops a tree repeatedly and then crafts a table, the video should be split at the moment the agent mines the last block of wood. Further details are provided in Section B.3.
Method [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average
VPT original 11.0% 29.0% 0.0% 22.0% 7.0% 23.0% 77.0% 14.0% 73.0% 33.0% 30.0% 38.0% -
GROOT original 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 -
ours 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 -
$\Delta$ +42.9% +107.7% 0 -9.3% +30.1% +166.7% +23.3% +50.0% +18.4% +131.6% +36.1% +166.3% +63.7%
STEVE-1 original 0.0% 1.0% 33.3% 33.3% 18.8% 65.6% 96.9% 18.8% 80.2% 57.3% 44.8% 46.9% -
ours 0.0% 3.1% 40.6% 77.1% 42.7% 71.9% 96.9% 47.9% 84.4% 67.7% 43.8% 71.9% -
$\Delta$ 0 -- +21.9% +131.3% +127.8% +9.5% 0 +155.6% +5.2% +18.2% -2.3% +53.3% +52.1%
Table 2: Success rate of different policies on Minecraft skill benchmarks. For VPT (Baker et al., 2022), we report the results of the behavioral cloning version. For GROOT (Cai et al., 2023b) and STEVE-1 (Lifshitz et al., 2023), we report the results of original and our re-trained with SBD, respectively. A value with % indicates the average success rate, while a value without % indicates the average rewards. The seeds for the Minecraft environment are fixed for the corresponding task to make a fair comparison between different models. Details of the tasks are provided in Section B.1.
Policies for Atomic Skills
We employ both language-conditioned and video-conditioned policies to learn atomic skills from the segmented demonstration videos as shown in Stage III. The video-conditioned policy is built upon GROOT (Cai et al., 2023b), which utilizes self-supervised learning to model instructions and behaviors. In the original GROOT, a fixed-length sequence of 128 frames serves as the instruction, which is encoded into a goal embedding using a conditioned variational autoencoder (C-VAE) (Kingma and Welling, 2014; Sohn et al., 2015). The decoder then predicts actions based on the instruction and environmental observations in an auto-regressive manner. The language-conditioned policy is derived from STEVE-1 (Lifshitz et al., 2023), an instruction-tuned VPT model capable of following open-ended texts or 16-frame visual instructions. This model is trained by first adapting the pretrained VPT model to follow commands in MineCLIP’s (Fan et al., 2022) latent space, and then training a C-VAE model to predict latent codes from the text. The original GROOT and STEVE-1 policies divide the training trajectories into segments of 128 frames and uniformly sample between 15 and 200 frames, respectively. We re-train these models using our SBD method on the segmented trajectories and compare the results to the original models to demonstrate the effectiveness of our approach. We retain the minimum and maximum length settings. The algorithm for pruning the length of each trajectory is detailed in B.4. Most of the original hyperparameter settings are retained as specified in the original papers. A complete list of modified hyperparameters is provided in Section B.2.
Method Wood Food Stone Iron Average
original 95% 44% 82% 32% -
ours 96% 55% 90% 35% -
$\Delta$ +1.1% +25.0% +9.8% +9.4% +11.3%
(a) OmniJARVIS (behavior cloning)
Method Wood Oak Birch Stone Iron Diamond Armor Food Average
original 92% 89% 90% 90% 33% 8% 12% 39% -
ours 97% 95% 94% 91% 35% 10% 19% 62% -
$\Delta$ +5.4% +6.7% +4.4% +1.1% +6.1% +25.0% +58.3% +59.0% +20.8%
(b) JARVIS-1 (in-context learning)
Table 3: Success rate of two agents with their corresponding controllers trained on dataset segmented by our SBD method and the original sequential sampling (SS) method on groups of long programmatic tasks. All tasks are tested 30 times. In each group, the agent is required to obtain a certain type of items from scratch or given an iron pickaxe. For example, the diamond group includes diamond pickaxe, diamond sword, jukebox, etc.
Task [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average
- 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 64.2
Info 33.0% 52.0% 100.0% 88.0% 97.0% 32.0% 23.7 65.0% 92.0% 47.0% 15.5 21.6 87.4
Loss 44.0% 54.0% 100.0% 94.0% 72.0% 46.0% 25.8 63.0% 95.0% 48.0% 17.9 22.7 91.8
Both 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 93.3
Table 4: Ablation on the components within SBD. We report the evaluation results on Minecraft atomic skills from the sequential sampling (- in the table) and SBD with different components.
Long-term Instruction Following
We use the hierarchical agents to model the long trajectories, which are widely used by (Driess et al., 2023; Wang et al., 2025, 2023b). We utilize in-context learning, as demonstrated by JARVIS-1 (Wang et al., 2024b), and imitation learning, as shown by OmniJARVIS (Wang et al., 2025), to evaluate the long-horizon instruction-following capabilities of hierarchical agents. JARVIS-1 is a hierarchical agent that uses the text-conditioned STEVE-1 as its policy and pretrained vision-language models as planners. We replace the text-conditioned policy with our re-trained STEVE-1 based on SBD. OmniJARVIS is a vision-language action agent built on FSQ-GROOT, which encodes instructions into discrete tokens rather than continuous embeddings. It is trained on behavior trajectories encoded into unified token sequences.
4 Experiments and Analysis
In our experiments, we answer the following questions:
$\bullet$ The short segments obtained through SBD method, are they more consistent at the semantic level, and can a better atomic skill policy be trained on these segments?
$\bullet$ Does using short videos segmented with new methods as skills perform better on long-horizon tasks?
$\bullet$ Which component of SBD is the most important?
4.1 Experimental Setups
Environmental Setups and Datasets
We take the Minecraft environment (Guss et al., 2019; Zheng et al., 2023) as the evaluation simulator. Within the Minecraft environment, we use OpenAI’s contractor dataset 7.x (early game) (Baker et al., 2022) as trajectories. The dataset contains offline trajectories with 68M frames with a duration of approximately 1000 hours, with at least half of the data from the first 30 minutes of the game. Our method generates a segmented dataset of 130k sub-trajectories using bc-early-game-3x (Baker et al., 2022) as the pretrained unconditional model.
Evaluation Benchmarks
For the atomic conditioned policies, we selected basic skills such as chop down trees and advanced skills like smelt items with furnace in Minecraft as evaluation benchmarks. We tested 12 different skill sets designed in MCU (Lin et al., 2023). Each task was tested over 100 times, except for sleep in bed and use bow, which were evaluated 10 times using the human Elo rating. The evaluation metrics are computed with the success rates and rewards. The hierarchical agents are evaluated on long-horizon programmatic tasks requiring the agents to start from an empty inventory in a new world until obtaining the final required items, such as obtain an iron pickaxe from scratch, which is usually a chain of atomic tasks. We split the long-horizon tasks into different groups, including wooden items, food, stones items, iron etc. For each group, we choose the tasks from JARVIS-1 benchmarks (Wang et al., 2024b) and evaluate each task over 10 times.
4.2 Main Results
In this section, we present the results of our method, SBD, applied to the two policies and two agents mentioned earlier. SBD achieves an average performance increase of 63.7% for GROOT and 52.1% for STEVE-1 compared to the original methods. Additionally, it enhances the performance of OmniJARVIS and JARVIS-1 by 11.3% and 20.8%.
Short-horizon Atomic Tasks
The results are shown in Table 2. Scores with % indicate success rates, while those without % represent rewards for the corresponding tasks. The controllers showed substantial improvements across most tasks, with an average performance enhancement of 63.7% for video-conditioned policy and 52.1% for language-conditioned policy.
Long-horizon Programmatic Tasks
To further verify that the improvements in controllers enhance the ability of hierarchical agents, we evaluate OmniJARVIS and JARVIS-1 on programmatic tasks selected from the original papers (Wang et al., 2025, 2024b). As shown in Table 3, the agents have substantial improvements across most of the tasks, achieving an average performance enhancement of 11.3% for Omnijarvis and 20.8% for JARVIS-1.
4.3 Ablation Study
(a)    Info only.
(b)    Loss only.
(c)    Both.
Figure 2: The length distribution of segments, split by info and loss. The info-only method is intrinsically semantically meaningful, suggesting that the loss-only method also identifies a semantically meaningful segmentation pattern. Furthermore, the similarity between the combined method and the loss-only method indicates that predictive loss is the primary factor in learning the segmentation pattern.
Skill Videos
Figure 3: Video Segment Examples. Top: sleep in bed. Bottom: collect grass. Each segment is accompanied by five screenshots. The first and last screenshots represent the initial and final frames of the segment, respectively. The remaining three screenshots are manually selected to best illustrate the skill’s progression. More segments can be found in Appendix C.
As introduced in Sections 3.2.1 and 3.2.2, the environment information and the skill boundary detection loss are important components for accurate segmentation of long videos. In this section, we explore the effectiveness of every component. So we compare the full SBD with sequential sampling (w/o loss and info), SBD w/ info only, and SBD w/ loss only. To evaluate the effectiveness of these configurations, we conducted an ablation study on the performance of four methods under GROOT. The results are presented in Table 4. Our findings are as follows: (i) Using external information alone significantly improves performance, with an average increase of 36.1% over the original agent. (ii) Using loss alone results in a better performance (an average increase of 5.0%) compared to using external information alone, highlighting the effectiveness of the core component of our method. This demonstrates that our method achieves strong performance even without external information. (iii) Combining loss and external information yields the best overall performance. This combination outperforms the individual components in half of the tasks, suggesting that each component captures unique segmentation patterns that the other misses.
We observe performance declines when employ both components in tasks smelt food and collect seagrass. This may be due to both components partitioning the trajectory at the same intervals but not at identical steps, leading to redundant splits and incorrect trajectory fragments.
4.4 Visualizations
Length Distribution
Fig. 2 shows the length distribution of sub-trajectories split by different methods with or without the two components of SBD. We observe that both loss-only and info-only methods follow similar distributions, where the log length and frequency of sub-trajectories exhibit a normal distribution. Since the info-only method is intrinsically semantically meaningful, this provides indirect evidence that the loss-only method also identifies a semantically meaningful segmentation pattern as the info-only method. Furthermore, when combining loss with external information, the distribution does not significantly differ from the loss-only method. This observation highlights that the predictive loss is the key driver in learning the segmentation pattern, with the external information serving a more supplementary role.
We sample one long unsegmented video segmented by our method for each skill in the datasets, such as sleep in bed and collect seagrass (see Fig. 3). More segments can be seen in Appendix C.
5 Related Works
Learning from Unsegmented Demonstrations
Prior work on learning from unsegmented demonstrations has proposed various methods to segment trajectories into sub-trajectories. For instance, some methods (Shankar and Gupta, 2020; Tanneberg et al., 2021; Kipf et al., 2019) utilize a variational autoencoder (Kingma and Welling, 2014) model to generate pairs of skill types and durations, $(g_{i},t_{i})_{i=1}^{k}$, from a given trajectory, thereby enabling segmentation. TACO (Shiarlis et al., 2018) adopts a weakly supervised framework that identifies task boundaries, $(t_{1},\ldots t_{k})$ based on task sketches, $(g_{1},\ldots g_{k})$, by solving a sequence alignment problem to achieve appropriate segmentation. BUDS (Zhu et al., 2022) constructs hierarchical task structures of demonstration sequences using a bottom-up strategy, deriving temporal segments through agglomerative clustering of the actions. FAST (Pertsch et al., 2025) proposes an approach based on byte-pair encoding (Gage, 1994) algorithm for segmentation and tokenization of robot action trajectories via time-series compression.
Option Learning
Option learning (Sutton et al., 1999) is a framework designed to learn temporal abstractions (i.e., skills or primitives) directly from the environment’s reward signal, primarily through online reinforcement learning. Recent approaches (Bacon et al., 2017; Bagaria and Konidaris, 2019; Klissarov and Precup, 2021) rely on policy gradient methods to learn the option policies. This framework requires reward signals, and DIAYN (Eysenbach et al., 2019) proposes a method to generate a reward function adaptively in the absence of reward signals, resulting in the unsupervised emergence of diverse skills.
Event Segmentation Theory
In neuroscience, episodic memory is the memory of everyday events consisting of short slices of experience (Conway, 2009). Event Segmentation Theory (EST) (Zacks et al., 2007) offers a theoretical perspective on how the neurocognitive system splits a long flow of memory into short events. According to EST, observers build event models of the current situation to generate predictions of future perceptual input. When errors in predictions arise, an event boundary is perceived, causing the event model to be reset and rebuilt. Our method employs a pretrained model to predict the agent’s action based on past observations and uses the predictive loss as an indicator to detect skill boundaries.
Agents in Minecraft
Many agents have been developed to interact with Minecraft environments (Zhao et al., 2024a; Jiang et al., 2025). For short-horizon tasks, methods typically employ imitation or reinforcement learning, as seen in works like VPT (Baker et al., 2022), which annotated a large YouTube Minecraft video dataset with actions and trained the first foundation agent in the domain using behavior cloning, and its derivatives (Lifshitz et al., 2023; Cai et al., 2023b; Yuan et al., 2024; Cai et al., 2023a, 2024b). For long-horizon, programmatic tasks, large language models (LLMs) are used as planners combined with skill policies (Wang et al., 2023b; Zhou et al., 2024; Yuan et al., 2023; Qin et al., 2024; Wang et al., 2024b; Li et al., 2024), and some methods (Wang et al., 2024b; Zhu et al., 2023; Wang et al., 2024a) employ explicit memory mechanisms to further enhance the long-horizon capabilities of LLM agents. Additionally, recent advances (Wang et al., 2025; Zhao et al., 2024b) have explored using end-to-end visual-language models (VLMs) to directly follow human instructions. Finally, some research (Guo et al., 2024; Zhang et al., 2023) focuses on open-ended creative tasks such as building and decoration, which often cannot be directly defined by rule-based rewards.
6 Conclusions
In this paper, we propose a novel temporal video segmentation algorithm to address the challenges of open-world skill discovery from unsegmented demonstrations. To generate a segmented dataset of video clips with independent skills, our algorithm detects the potential skill boundaries based on the predictive loss of a pretrained action-prediction model. External boundary indicators can also be incorporated in our method. The method does not require any manual annotations and can be employed directly on massive internet gameplay videos.
7 Limitations and Future work
The current algorithm is relatively time-consuming, requiring approximately 1 minute of processing on an NVIDIA RTX 4090Ti for every 5-minute video, due to the need to predict actions and compute loss across the entire trajectory. As a result, we have not tested our method on larger datasets. Moreover, evaluating the method requires reprocessing the entire dataset and retraining the controllers, so we have not conducted an extensive ablation study on pretrained model selection or hyperparameter tuning. Additionally, we observe that the algorithm becomes unstable in action-intensive scenarios (e.g., combat situations), often generating many short segments. Future work could focus on improving the efficiency and stability of the algorithm, as well as exploring more effective evaluation methods. We also anticipate the potential application of our method to larger datasets and other open-world video games beyond Minecraft.
Acknowledgement
This work is funded in part by the National Science and Technology Major Project 2022ZD0114902. We thank a grant from CCF-Baidu Open Fund.
Appendix A Proofs
In this section, we provide a detailed proof of Theorem 3.4, regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition. We prove the lower bound and upper bound respectively in the following two subsections.
A.1 Proof of Upper Bound Under Skill Non-Transition
$P(\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}>\frac{(K-1)c}{K})$ $>P(P(a_{t+1}|o_{1:t+1})>\frac{(K-1)c}{K})$
[…]
$>P(\pi_{t}(a_{t+1}|o_{1:t+1})>c)$ (Assumption 3.1)
$=P(\pi_{t+1}(a_{t+1}|o_{1:t+1})>c)$
$>1-\delta$ (Assumption 3.2)
A.2 Proof of Lower Bound Under Skill Transition
We first magnify the likelihood ratio between the next action and the average action history,
$\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}$ $<\frac{P(\pi_{t+1}=\pi_{t}|o_{1:t+1})\pi_{t}(a_{t+1}|o_{1:t+1})+P(\pi_{t+1}\neq\pi|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}$ (LABEL:eq:transition)
$<\frac{P(\pi_{t+1}=\pi_{t}|o_{1:t+1})\pi_{t}(a_{t+1}|o_{1:t+1})+P(\pi_{t+1}\neq\pi|o_{1:i})}{\frac{K-1}{K}(\prod_{i=1}^{t}\pi_{i-1}(a_{i}|o_{1:i}))^{1/t}}$ (LABEL:eq:transition and
Assumption 3.1)
$<\frac{\pi_{t}(a_{t+1}|o_{1:t+1})+\frac{1}{K}}{\frac{K-1}{K}(\prod_{i=1}^{t}\pi_{i-1}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption 3.1)
$=\frac{K}{K-1}\frac{\pi_{t}(a_{t+1}|o_{1:t+1})}{\prod_{i=1}^{t}\pi_{t}(a_{i}|o_{1:i}))^{1/t}}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}$
$<\frac{Km}{2(K-1)}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption 3.3)
Therefore,
$P(\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}<\frac{Km}{2(K-1)}+\frac{1}{c(K-1)})$
$>P(\frac{Km}{2(K-1)}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}<\frac{Km}{2(K-1)}+\frac{1}{c(K-1)})$
$=P(\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}>c)$
$>\prod_{i=1}^{t}P(\pi_{i}(a_{i}|o_{1:i})>c)$
$>(1-\delta)^{t}>1-t\delta$ (Assumption 3.2)
Appendix B Experiment Details
In this section, we provide our experiment details including benchmark, training details for GROOT and STEVE-1, how we use event-based information as external auxiliary information, and the length pruning algorithm for sub-trajectories.
B.1 Tasks in the Early Game Benchmark
Minecraft SkillForge is a diverse benchmark of 30 short atomic tasks that can comprehensively evaluate the mastery of atomic skills by agents in Minecraft. We choose 10 early game tasks from the original benchmark which the original agent can already handle at a basic level because our method is an improvement on dataset processing instead of a new model architecture, so it cannot help agents learn new skills that they are completely incapable of acquiring previously. Besides, we add the task "use torch", which is also a useful skill in Minecraft; and we add an enhanced version of the task "collect wood", which requires the agent to start from the plains instead of the forest, aiming to test its ability to explore and find trees.
Details of the 12 tasks in our early game benchmark are shown in Table 5. For each task, we include the evaluation metric and a brief description of what the task is. For tasks "Sleep" and "Use bow", GROOT performs very well on the original metric, so we design a manual metric that better assesses its actual performance. STEVE-1 can take text as prompts, which are also listed in the table. For some of the tasks, it is tricky to find proper text prompts, so we use visual prompts instead.
Task Metric Description Text Prompt for STEVE-1
[Uncaptioned image] Craft item cooked mutton Given a furnace and some mutton and coal, craft a cooked mutton. -
[Uncaptioned image] Kill entity sheep Summon some sheep before the agent, hunt the sheep. -
[Uncaptioned image] STEVE-1: Use item white bed. GROOT: Sleep in bed properly. Given a white bed, sleep on it. Sleep in bed.
[Uncaptioned image] Use item torch Give some torches, use them to light up an area. Time is set at night. Use a torch to light up an area.
[Uncaptioned image] Use item birch boat Given a birch boat, use it to travel on water. Biome is ocean. -
[Uncaptioned image] STEVE-1: Use item bow. GROOT: Use item bow 20%, Shoot in distance 40%, take aim 40% Given a bow and some arrows, shoot the sheep summoned before the agent. -
[Uncaptioned image] Mine block stone (cobblestone, iron, coal, diamond) Given an iron pickaxe, collect stone starting from cave. Night vision is enabled. Collect stone.
[Uncaptioned image] Mine block seagrass (tall seagrass, kelp) Given an iron pickaxe, collect seagrass starting from ocean. -
[Uncaptioned image] Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, collect wood starting from forest. Chop a tree.
[Uncaptioned image] Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, find and collect wood starting from plains. Chop a tree.
[Uncaptioned image] Mine block dirt (grass block) Given an iron pickaxe, collect dirt starting from plains. Collect dirt.
[Uncaptioned image] Mine block grass (tall grass) Given an iron pickaxe, collect grass starting from plains. Collect grass.
Table 5: Details of 12 atomic tasks in our early game benchmark for testing GROOT and STEVE-1.
B.2 Training Details
The modified hyperparameters for GROOT and STEVE-1 are listed in Table 6. We adjust parallel GPUs, gradient accumulation batches, and batch sizes to better align with our available computing resources. To speed up convergence without compromising performance, we double the learning rate for GROOT. The total number of frames for STEVE-1 is also modified, as we change the training dataset from a mixed subset of 8.x (house building from scratch), 9.x (house building from random materials), and 10.x (obtaining a diamond pickaxe) to the full 7.x dataset (early game), which better suits our benchmark tasks.  [^1] OpenAI released five subsets of contractor data: 6.x, 7.x, 8.x, 9.x, and 10.x.
All models are trained parallelly on four NVIDIA RTX 4090Ti GPUs. We follow the same policy training pipeline as the original paper except for the modified hyperparameters mentioned above. GROOT is trained under three epochs of our dataset. STEVE-1, however, is only trained under 1.5 epochs of our dataset, because we observe that the model starts to overfit on the dataset if it is trained further.
Hyperparameter Value
Learning Rate 0.00004
Parallel GPUs 4
Accumulate Gradient Batches 1
Batch Size 8
Hyperparameter Value
Parallel GPUs 4
Accumulate Gradient Batches 4
Batch Size 4
n_frames 100M
Table 6: Modified hyperparameters for training controllers. (Left) GROOT. (Right) STEVE-1.
B.3 Event-Based Information
Within the Minecraft environment, we focus on events in the categories "use item", "mine block", "craft item", and "kill entity", as they reflect the player’s primary activities. Multiple events may occur simultaneously and are recorded as a set. Only the final step of a repeating sequence of event sets is marked as positive. For example, in the sequence: ("use item iron pickaxe", "mine block iron ore"), ("use item iron pickaxe", "mine block iron ore"), ("use item iron pickaxe", "mine block diamond ore"), ("use item torch"), ("use item torch"), the second, third, fifth steps will be marked as positive. Additionally, for any step $t$ that includes a "kill entity" event, we mark $t+16$ as positive instead of $t$, because there will be a short death animation that follows the entity’s death and we want it to be included in the same segment.
B.4 Length Pruning Algorithm
STEVE-1 forces a minimum and maximum length of trajectories in its method so we need a length pruning algorithm. In our implementation, the length of each trajectory is pruned as follows: If a trajectory is too short, it is merged with subsequent trajectories until it meets the minimum length requirement. If this results in a trajectory exceeding the maximum length, it is truncated, and the remainder forms the beginning of the next trajectory. For instance, given a minimum and maximum length of 15 and 200, sub-trajectories of lengths 12, 12, 6, 196, and 37 are adjusted to 24, 200, and 39. Here, the first two sub-trajectories are merged, while the 6 is combined with 196 but truncated at 200, with the remaining 2 merged into the next trajectory. There might be more sophisticated strategies, but we use this straightforward algorithm for simplicity.
Appendix C Examples of Skill Videos
We sample one video segmented by our method for each skill in the benchmark, presenting each video with five screenshots. The first and last screenshots correspond to the first and last frames of the video, while the other three are manually selected to best illustrate the progression of the skill.
  • smelt food
  • hunt sheep
  • sleep
  • use torch
  • use boat
  • use bow
  • collect stone
  • collect seagrass
  • collect wood
  • collect dirt
  • collect grass

### v2

@@
-Open-World Skill Discovery from Unsegmented Demonstrations
+Open-World Skill Discovery from Unsegmented Demonstration Videos
 Abstract
@@
 Online demonstration videos are typically long but unsegmented, making them difficult to segment and label with skill identifiers.
-Unlike existing methods that rely on sequence sampling or human labeling, we have developed a self-supervised learning-based approach to segment these long videos into a series of semantic-aware and skill-consistent segments.
+Unlike existing methods that rely on random splitting or human labeling, we have developed a self-supervised learning-based approach to segment these long videos into a series of semantic-aware and skill-consistent segments.
 Drawing inspiration from human cognitive event segmentation theory, we introduce Skill Boundary Detection (SBD), an annotation-free temporal video segmentation algorithm.
@@
 We evaluated our method in Minecraft, a rich open-world simulator with extensive gameplay videos available online.
-Our SBD-generated segments improved the average performance of conditioned policies by 63.7% and 52.1% on short-term atomic skill tasks, and their corresponding hierarchical agents by 11.3% and 20.8% on long-horizon tasks.
+The SBD-generated segments yielded relative performance improvements of 63.7% and 52.1% for conditioned policies on short-term atomic tasks, and 11.3% and 20.8% for their corresponding hierarchical agents on long-horizon tasks, compared to random segmented baselines.
 Our method can leverage the diverse YouTube videos to train instruction-following agents.
-The project page can be found in https://craftjarvis.github.io/SkillDiscovery/.
-[…]
-Yitao Liang <yitaol@pku.edu.cn>
-[…]
-Jingwen Deng <dengjingwen@stu.pku.edu.cn>, Zihao Wang <zhwang@stu.pku.edu.cn>, Anji Liu <liuanji@cs.ucla.edu>
+The project page is at https://craftjarvis.github.io/SkillDiscovery/.
+footnotetext: Preprint version of the paper accepted at ICCV 2025.
 1 Introduction
 Method Illustration Type
-Sequential sampling [Uncaptioned image] rule-based
+Random splitting [Uncaptioned image] rule-based
 Reward-driven [Uncaptioned image] rule-based
@@
 Existing methods usually rely on human-designed rules, while our method is learning-based.
-Sequential sampling can result in a single skill spanning different segments or multiple skills located within one segment.
+Random splitting can result in a single skill spanning different segments or multiple skills located within one segment.
 Reward-driven methods require additional reward information, which is challenging for human annotators to label.
 Top-down methods often result in limited skill diversity and high computation costs.
-Bottom-up methods are limited in fully observable environments and hard in visually partially observable environments.
+Bottom-up methods struggle in visually partially observable environments.
 [Uncaptioned image] are visual observations and [Uncaptioned image] are actions.
-Most existing LLM-based instruction-following agents adopt a two-layer structure of planner and controller to complete tasks, which first convert the instruction into atom skills through the planner, and then use a conditioned policy to convert it into actions based on the current observation to interact with the environment (Wang et al., 2025, 2023b, 2024b; Li et al., 2024; Wang et al., 2024a; Cheng et al., 2024).
-When training these agents, we need to first convert long sequences into a series of short atom skill sequences, and then train the controller and planner separately.
-However, interactions in real-world videos are often long-term and unsegmented with detailed language skills.
+Most existing LLM-based instruction-following agents adopt a two-tier architecture of a planner and a controller.
+The planner decomposes high-level instructions into atomic skills, which the controller then maps to actions based on current observations for interaction with the environment [40, 38, 39, 26, 36, 9].
+Training such agents typically involves segmenting long trajectories into atomic skill sequences, followed by separate training of the planner and controller.
 Learning skills from long-sequence videos is critical for building such hierarchical agents.
-However, the concept of a “skill” is ill-defined and varies widely across domains such as video gaming (Cai et al., 2023b; Wang et al., 2024a; Hafner et al., 2023; Zhang et al., 2024), robotic control (Zitkovich et al., 2023), and autonomous driving (Wang et al., 2023a).
+However, real-world video data are often long-horizon, unsegmented, and with no fine-grained skill labels.
+In addition, the concept of a “skill” is ill-defined and varies widely across domains such as video gaming [6, 36, 19, 46], robotic control [53], and autonomous driving [37].
 This ambiguity, along with the immense diversity of skills in open worlds, makes skill learning particularly challenging, especially in partially observable settings.
-To enable open-world skill learning from unsegmented demonstrations (continuous streams of observation-action pairs without explicit labels), the first challenge is to segment these streams into semantically meaningful, self-contained skills.
-We list the existing segmentation methods in Table 1, including the plain sequential sampling, reward-driven methods, top-down, and bottom-up methods.
-The native sequential sampling segmentation methods (Cai et al., 2023b; Lifshitz et al., 2023), divide videos into segments of predefined lengths (e.g., fixed length, uniform distribution).
-However, these methods do not ensure that each segment contains a distinct skill.
-Additionally, predefined lengths may not match the actual distribution of skill lengths in real-world scenarios (see Section 4.4).
-Reward-driven methods (Sutton et al., 1999) discover skills through the environment’s reward signal.
-It is limited by its inability to capture skills that lack associated rewards and by the risk of splitting a single skill into multiple segments when rewards are repeatedly gained during execution.
-Top-down methods (Shiarlis et al., 2018; Belkhale et al., 2024) rely on predefined skill sets from human experts.
+To enable open-world skill learning from unsegmented demonstration videos, the first challenge is to segment them into semantically meaningful, self-contained skills.
+We list the existing segmentation methods in Tab.˜1.
+The naive random splitting methods [6, 27] divide videos into segments of predefined lengths (e.g., fixed length, uniform distribution).
+However, they do not ensure that each segment contains a distinct skill.
+Additionally, predefined lengths may not match the actual distribution of skill length in real world (see Sec.˜4.4).
+Reward-driven methods [34] discover skills through the environment’s reward signal.
+They cannot capture skills with no associated rewards and may split one skill into multiple segments when rewards are repeatedly gained during execution.
+Top-down methods [32, 4] rely on predefined skill sets from human experts.
 They use manual labeling or supervised learning to segment videos.
-Although this approach can produce reasonable results, it is expensive and limited by the narrow range of predefined skills.
-Bottom-up methods (Zhu et al., 2022; Pertsch et al., 2025) use algorithms such as agglomerative clustering or byte-pair encoding (BPE) (Gage, 1994) to split action sequences.
-However, they struggle in partially observable settings where both observations and actions must be considered because these algorithms cannot deal with observations due to the high dimensionality.
-All the above methods usually rely on human-designed rules to segment the unsegmented videos, which highlights the need for an effective and label-free method that can automatically segment skills from unsegmented demonstrations in open worlds.
-Inspired by event segmentation theories (EST) (Zacks et al., 2007), which propose that humans naturally partition continuous experiences into discrete events when prediction errors in perceptual expectations rise, we introduce Skill Boundary Detection (SBD)—a method for autonomously identifying potential skill boundaries within extended trajectories.
-SBD employs a predictive model trained on a dataset of unsegmented videos to forecast future actions based on past observations, effectively capturing temporal dependencies (Baker et al., 2022).
-We then use this pretrained model to make predictions on an unsegmented video and compare them to the ground truth actions.
-A significant increase in prediction error indicates a shift in the skill being executed (Theorem 3.4), enabling us to detect boundaries between different skills in the video.
+Although this method can produce reasonable results, it is expensive and limited by the narrow range of predefined skills.
+Bottom-up methods [52, 29] use algorithms such as agglomerative clustering or byte-pair encoding [16] to split action sequences.
+However, they struggle in visually partially observable settings where both observations and actions must be considered.
+All the methods above rely on human-designed rules to segment the videos, which highlights the need for a learning-based method that can adaptively segment skills from unsegmented videos in open world.
+Inspired by event segmentation theory (EST) [44], which proposes that humans naturally split continuous experiences into discrete events when prediction errors in perceptual expectations rise, we introduce Skill Boundary Detection (SBD)—a method for automatically identifying potential skill boundaries in long trajectories.
+SBD employs a predictive model trained on a dataset of unsegmented videos to predict future actions based on past observations, effectively capturing temporal dependencies [3].
+We then use this pretrained model to make predictions on unsegmented videos and compare them to the ground-truth actions.
+A significant increase in prediction error indicates a shift in the skill being executed (Theorem˜3.4), enabling us to detect boundaries between different skills.
 SBD relies on self-supervised learning, removing the need for additional human labeling.
 This allows SBD to utilize a wide range of YouTube videos to train instruction-following agents.
-We evaluate SBD in open-world Minecraft (Guss et al., 2019; Fan et al., 2022; Lin et al., 2023).
+We evaluate SBD in Minecraft [18, 15, 28], a rich open world simulator.
 First, we apply SBD to create a segmented Minecraft video dataset.
-We then train the video-conditioned policy (Cai et al., 2023b) and the language-conditioned policy (Lifshitz et al., 2023) on this dataset.
-We evaluate these policies on diverse Minecraft skills benchmark (Lin et al., 2023).
-The results show significant improvements compared to the existing policies (Cai et al., 2023b; Lifshitz et al., 2023), with performance increases of 63.7% and 52.1% compared to the original methods, respectively.
-We also test their corresponding hierarchical agents using behavior cloning and in-context learning, achieving performance increases of 11.3% and 20.8%, respectively.
-These findings underscore the effectiveness of our proposed method for skill discovery from unsegmented demonstrations and its potential to advance open-world skill learning.
+We then re-train a video-conditioned policy [6] and a language-conditioned policy [27] on this dataset.
+We evaluate them on a diverse Minecraft skills benchmark [28].
+The improved versions of these policies result in relative performance increases of 63.7% and 52.1% compared to the original ones, respectively.
+We also test their corresponding hierarchical agents that use behavioral cloning and in-context learning, achieving relative performance increases of 11.3% and 20.8%, respectively.
+These findings show the effectiveness of our proposed method for skill discovery from unsegmented demonstrations and its potential to advance open-world skill learning.
 Figure 1: Pipeline of our method SBD for discovering skills from unsegmented demonstration videos.
-Stage I: An unconditional Transformer-XL based policy model (Dai et al., 2019; Baker et al., 2022) is pretrained on an unsegmented dataset to predict future actions (labeled by an inverse dynamics model) based on past observations.
+Stage I: An unconditional Transformer-XL based policy model [11, 3] is pretrained on an unsegmented dataset to predict future actions (labeled by an inverse dynamics model) based on past observations using behavioral cloning.
 Stage II: The pretrained unconditional policy will produce a high predicted action loss when encountering uncertain observations (e.g., deciding whether to kill a new sheep) in open worlds.
@@
 Stage III: We train a conditional Transformer-XL based policy model on the segmented dataset to master a variety of atomic skills.
-Stage IV: Finally, we use hierarchical methods (a combination of vision-language models and the conditional policy) to model the long demonstration and follow long-horizon instructions.
+Stage IV: Finally, we use hierarchical methods (a combination of vision-language models and the conditional policy) to model the long demonstration videos and follow long-horizon instructions.
 2 Problem Formulation
 Skills in Behavioral Cloning
-Behavioral cloning learns a policy $\pi$ that maps observations $o$ from the observation space $\mathcal{O}$ to actions $a$ in the action space $\mathcal{A}$ using supervised learning on observation-action pairs $(o_{i},a_{i})_{i=1}^{T}$ obtained from expert demonstrations.
-In partially observable environments, past observations $o_{1:t}$ are often used instead of the current observation $o_{t}$ as input to the policy $\pi$.
-Additionally, to allow the agent to dynamically adjust its behavior based on its objective, recent advances (Khandelwal et al., 2021; Du et al., 2021; Xie et al., 2023) focus on learning goal-conditioned policies $\pi(o_{1:t},g)$, where $g$ represents the goal.
-This type of goal-conditioned policy can be considered a “skill”.
+Behavioral cloning learns a policy $\pi:\mathcal{O}\rightarrow\mathcal{A}$ from expert demonstrations, mapping observations to actions based on pairs $(o_{i},a_{i})_{i=1}^{T}$.
+In partially observable settings, the policy typically takes the observation history $o_{1:t}$ as input.
+To enable goal-directed behavior, recent works [22, 13, 41] learn goal-conditioned policies $\pi(o_{1:t},g)$, where $g$ denotes a target goal.
+Such goal-conditioned policies are referred to as skills, as they encode reusable behaviors for achieving specific goals.
 Hierarchical Agent Architecture
@@
 Goals such as build a house or mine diamond blocks involve numerous intermediate steps, making it computationally and conceptually difficult to model an agent’s behavior without breaking the goal into manageable parts.
-To tackle this, hierarchical agent architectures are often used (Wang et al., 2024a; Driess et al., 2023; Cai et al., 2024a; Iqbal et al., 2022).
+To tackle this, hierarchical agent architectures are often used [36, 12, 7, 20].
 These architectures consist of a high-level planner and a low-level controller.
@@
 where $\pi_{1}$ is the planner’s policy and $\pi_{2}$ is the controller’s policy.
-The planner periodically (not at every step) checks whether the current sub-goal has been finished and determines whether to change policy, so $g_{t}$ is relatively static and does not change frequently.
-Identifying implicit skills in offline trajectories
-A critical challenge in training hierarchical agents, as described above, is the identification of skills within the offline trajectory data.
+The planner periodically (not at every step) checks whether the current sub-goal has been finished and decides whether to change it, so $g_{t}$ does not change frequently.
+Identifying Implicit Skills in Trajectories
+A critical challenge in training hierarchical agents, as described above, is the identification of skills within the trajectory data.
 To be more specific, given $\tau=(o_{i},a_{i})_{i=1}^{T}$, we want to split it into $\tau_{1}=(o_{i},a_{i})_{i=1}^{t_{1}},\tau_{2}=(o_{i},a_{i})_{i=t_{1}}^{t_{2}},\ldots,\tau_{k}=(o_{i},a_{i})_{i=t_{k-1}}^{T}$ where each sub-trajectories $\tau_{j}$ shows an independent skill $\pi(g_{j})$.
-Since some controllers use self-supervised methods to learn sub-goals (Lifshitz et al., 2023; Cai et al., 2023b), we can ignore the language labels of goal $g_{j}$, but must identify $t_{j}$ (i.e., the moments at which the agent adopts a different skill).
+Since some controllers use self-supervised methods to learn sub-goals [27, 6], we don’t have to identify $g_{j}$, but must identify $t_{j}$ (i.e., the moments at which the agent adopts a different skill).
 For clarity, in the subsequent sections, we denote the skill adopted at a certain time step $t$ as $\pi_{t}$, corresponding to $\pi(g_{t})$.
@@
 We first introduce our overall pipeline to build policies and long-horizon hierarchical agents based on unsegmented videos in Section 3.1.
-Then we introduce how we split the unsegmented demonstration into our policy and agent learning sequences in Section 3.2.
-We finally introduce the implementation details on Section 3.3.
+Then we introduce how we split the unsegmented videos into skill sequences in Section 3.2.
+We finally introduce the implementation details in Section 3.3.
 3.1 Pipeline
-Given a long sequence of trajectories $D={(o_{i},a_{i})}_{i=1}^{T}$ with interleaved visual observations $o_{i}$ and actions $a_{i}$, we first train an unconditioned policy $\pi_{\text{uncondition}}$ using imitation learning to predict the actions as follows:
-$\min_{\theta}\sum_{t\in[1\dots T]}-\log\pi_{\text{uncondition}}(a_{t}|o_{1},\dots,o_{t}).$ (2)
-When training on action label-free YouTube videos, the action labels are generated from Inverse Dynamics Model (Baker et al., 2022).
-We then evaluate this unconditioned policy $\pi_{\text{uncondtion}}$ on the entire dataset $D$ of long sequences.
-We developed an automated method called SBD to segment the complete video into a series of short segments $D_{seg}={(o_{i},a_{i})}_{i=m}^{n}$ based on these evaluation results, where $m$ and $n$ is the selected segmentation timestamps.
-These short segments $D_{seg}$ are then used for skill learning to obtain conditioned policies (Cai et al., 2023b; Lifshitz et al., 2023).
-Finally, we integrate these skills with vision language models to build hierarchical agents, enabling them to learn and follow long-horizon tasks in long video sequences (Wang et al., 2025, 2024b).
-We show the overall pipeline is illuminated in Figure 1.
-3.2 Skill Boundary Detection
+Given a long sequence of trajectories $D={(o_{i},a_{i})}_{i=1}^{T}$, we first train an unconditional policy $\pi_{\text{unconditional}}$ using behavioral cloning to predict the actions purely based on past observations as follows:
+$\min_{\theta}\sum_{t\in[1\dots T]}-\log\pi_{\text{unconditional}}(a_{t}|o_{1:t}).$ (2)
+When training on action label-free videos, the action labels are generated from an Inverse Dynamics Model [3] .
+We then evaluate this unconditional policy $\pi_{\text{uncondtional}}$ on the entire dataset $D$ of long sequences.
+Our proposed method SBD segments the complete video into a series of short segments $D_{seg}={(o_{i},a_{i})}_{i=m}^{n}$ based on the evaluation results, where $m$ and $n$ is the selected segmentation timestamps.
+$D_{seg}$ is then used for skill learning to obtain conditional policies [6, 27] that take additional videos or text as instructions.
+Finally, we integrate these skills with vision language models to build hierarchical agents, enabling them to complete long-horizon tasks [40, 39].
+See the overall pipeline in Fig.˜1.
+3.2 Skill Boundary Detection (SBD)
 Algorithm 1 Skill Boundary Detection
 1:Input: $(o_{i},a_{i},e_{i})_{i=1}^{T}$, a model $M$ to predict $a_{t}$ given $o_{1:t}$. $e_{i}$ is the boolean external information indicating whether this step should be a boundary.
-[…]
+2:Initialize: $\text{begin}\leftarrow 1$, $\text{loss\_history}\leftarrow[]$, $\text{boundaries}\leftarrow[]$
 3:for $t\leftarrow 1$ to $T$ do
 4:  $\text{loss}\leftarrow M(a_{t}\mid o_{\text{begin}:t})$
-[…]
-[…]
-[…]
+5:  $\text{loss\_history.append}(\text{loss})$
+6:  if $\text{loss}-\text{mean}(\text{loss\_history})>\text{GAP}$ or $e_{t}$ is true then
+7:   $\text{boundaries.append}(t)$
 8:   $\text{begin}\leftarrow t$
-[…]
+9:   $\text{loss\_history}\leftarrow[]$
 10:  end if
@@
 12:Return: boundaries
-SBD takes a long unsegmented trajectory as input and learns an unconditional model to predict the next action based on previous observations.
-The model does not know about implicit skill transitions.
-To simulate the model’s memory, a sliding window is used.
-At each time step $t$, the model predicts the next action and compares it with the ground truth to compute the loss.
-A skill transition is considered likely if the loss exceeds the average loss by a hyperparameter, GAP, or if an external indicator is true.
+SBD takes a long unsegmented trajectory and the unconditional action-prediction model as input.
+A sliding window is used to simulate the model’s memory of past observations.
+At each time step $t$, we use the model to predict the next action and compare it with the ground truth to compute the loss.
+A skill change is considered likely if the loss exceeds the average loss by a hyperparameter, GAP, or if an external indicator is true.
 In such cases, $t$ is marked as a boundary.
 The model’s memory is then cleared, and the algorithm proceeds to analyze the next sub-trajectory.
-The two core components of the algorithm are loss and external information.
+We then explain the two core components of the algorithm: loss and external information.
 3.2.1 Boundary with Entropy Loss
 In this section, we explain the core idea of our method: why loss can indicate skill boundaries.
+Intuitively, without goal information, similar observations in an open world can inherently lead to different actions—even with perfect modeling.
+For example, upon seeing a sheep, both ignore and attack may be plausible depending on the (unknown) goal.
+Skill transitions are moments where such ambiguity is highest.
+Since continuing a skill is more likely than changing it, the policy learns to favor the former, making skill transitions statistically surprising.
+According to Bayes’ rule, when the loss of the policy increases largely, it is likely that a skill transition has occurred.
 To support this, we introduce three key assumptions about skills: skill consistency, skill confidence, and action deviance at skill transition.
-The first assumption is that the skill being used does not change frequently.
-The parameter $K$ is a large number that determines how rare these changes are.
-The idea is that the agent should consistently stick to a skill unless there is a strong reason to switch to another, which would result in an increase in predictive loss for detecting a skill transition.
+The first assumption is that the skill being used does not change frequently (less than $1/K$).
+The idea is that the agent should consistently stick to a skill unless there is a strong reason to change to another, which would result in an increase in predictive loss.
 Assumption 3.1 (Skill Consistency).
@@
 The second assumption states that, at any given time, the agent is confident in the action it takes.
-This means the probability of the agent choosing an action with high confidence is very high.
-The parameter $c$ sets the minimum confidence level, while $\delta$ is a small value representing the frequency with which the policy might act with lower confidence.
+In other words, the probability of the agent choosing an action with confidence $c$ is greater than $1-\delta$.
 This assumption ensures that the policy reliably makes decisions based on its learned skills, rather than being uncertain in its actions.
@@
 This is because, in LABEL:eq:transition, in the first term the probability of $a_{t+1}$ under the original skill $\pi_{t}(a_{t+1}|o_{1:t+1})$ is low, and in the second term the probability of skill transition is low.
-On the contrary, if the agent does not change its skill, both terms will be high so the predictive probability will be high.
+On the contrary, if the agent does not change its skill, both terms will be high, so the predictive probability will be high.
 Therefore, we have the following theorem regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
@@
 Proof.
-See Appendix A for the detailed proof. ∎
-If $c>m$ and $(K-4)c^{2}>2$, then the lower bound of relative predictive probability under skill transition in Theorem 3.4 is greater than the upper bound of it under skill non-transition.
+See Appendix˜B for the detailed proof. ∎
+If $c>m$ and $(K-4)c^{2}>2$, then the lower bound of relative predictive probability under skill transition in Theorem˜3.4 is greater than the upper bound of it under skill non-transition.
 Therefore, the theorem demonstrates that the model predicts actions less accurately when the agent changes its skill.
-Our action-prediction model uses the negative log-likelihood of actions $-\log P(a_{t}|o_{1:t})$ as loss.
-Consequently, at line 6 of Algorithm 1, when the loss exceeds the average loss by a predefined threshold, it is likely that a skill transition has occurred.
-3.2.2 Boundary with Auxiliary Information
-In Section 3.2.1, we demonstrate that our algorithm can identify the first distinguishable skill boundary using only observations and actions.
-However, some datasets include additional external information, such as privileged in-game data or VLM detections.
-Incorporating these auxiliary signals improves the detection of skill boundaries that are otherwise difficult to recognize through loss-based boundary detection alone.
-For example, in Minecraft, crafting an item does not involve abrupt changes in mouse movement, making it hard to detect boundaries purely from predictive action loss.
+Our action-prediction model uses the negative log-likelihood of actions $\log P(a_{t}|o_{1:t})$ as loss.
+Consequently, at line 6 of Algorithm˜1, when the loss exceeds the average loss by a predefined threshold, it is likely that a skill transition has occurred.
+3.2.2 Boundary with External Information
+In Sec.˜3.2.1, we demonstrate that SBD can identify the first distinguishable skill boundary using only observations and actions.
+However, some datasets include additional external information such as privileged in-game data or VLM annotations [guo2025seed1, comanici2025gemini].
+Incorporating these auxiliary signals improves the detection of skill boundaries that are otherwise difficult to recognize through loss-based detection alone.
+For example, in Minecraft, crafting an item does not involve abrupt changes in mouse movement, making it hard to detect boundaries purely with predictive action loss.
 In such cases, in-game logs of successful crafting events provide valuable hints.
-It is important to note that this is an optional component of our algorithm; as we will discuss in Section 4.3, it also performs well on datasets without external auxiliary information.
-3.3 Implemetation Details
-Unconditioned Policy and Segmentation Details
-For the unconditioned policy $\pi_{\text{uncondition}}$ in Stage I, we use the pre-trained vpt-3x models (Baker et al., 2022).
-This is a foundational Transformer-XL (Dai et al., 2019) based model pretrained on large-scale YouTube data and finetuned on the early-game dataset using behavioral cloning.
-The hyperparameter GAP in Stage II is set to 18 considering the average trajectory length and semantics.
-As for the external information, we use event-based information such as mine_block:oak_log or use_item:torch available in the contractor dataset (Baker et al., 2022).
-To identify potential skill transitions, only the final event in a series of identical consecutive events is marked as positive, as it signifies a possible skill transition.
-For instance, if the agent chops a tree repeatedly and then crafts a table, the video should be split at the moment the agent mines the last block of wood.
-Further details are provided in Section B.3.
-Method [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average
+It is important to note that this is an optional component of SBD; as we will discuss in Sec.˜4.3, it also performs well on datasets without external auxiliary information.
+Task
+use
+furnace
+hunt
+sheep
+sleep
+in bed
+use
+torch
+use
+boat
+use
+bow
+collect
+stone
+collect
+seagrass
+collect
+wood
+find and
+collect wood
+collect
+dirt
+collect
+grass
+Average
 VPT original 11.0% 29.0% 0.0% 22.0% 7.0% 23.0% 77.0% 14.0% 73.0% 33.0% 30.0% 38.0% -
-GROOT original 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 -
+GROOT
+(video
+conditioned)
+original 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 -
 ours 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 -
 $\Delta$ +42.9% +107.7% 0 -9.3% +30.1% +166.7% +23.3% +50.0% +18.4% +131.6% +36.1% +166.3% +63.7%
-STEVE-1 original 0.0% 1.0% 33.3% 33.3% 18.8% 65.6% 96.9% 18.8% 80.2% 57.3% 44.8% 46.9% -
+STEVE-1
+(image & text
+conditioned
+original 0.0% 1.0% 33.3% 33.3% 18.8% 65.6% 96.9% 18.8% 80.2% 57.3% 44.8% 46.9% -
 ours 0.0% 3.1% 40.6% 77.1% 42.7% 71.9% 96.9% 47.9% 84.4% 67.7% 43.8% 71.9% -
-$\Delta$ 0 -- +21.9% +131.3% +127.8% +9.5% 0 +155.6% +5.2% +18.2% -2.3% +53.3% +52.1%
-Table 2: Success rate of different policies on Minecraft skill benchmarks.
-For VPT (Baker et al., 2022), we report the results of the behavioral cloning version.
-For GROOT (Cai et al., 2023b) and STEVE-1 (Lifshitz et al., 2023), we report the results of original and our re-trained with SBD, respectively.
+$\Delta$ 0 - +21.9% +131.3% +127.8% +9.5% 0 +155.6% +5.2% +18.2% -2.3% +53.3% +52.1%
+Table 2: Success rate of different policies on Minecraft skill benchmark.
+For VPT [3], we report the results of the behavioral cloning version.
+For GROOT [6] and STEVE-1 [27], we report the results of the original and our re-trained models with SBD, respectively.
 A value with % indicates the average success rate, while a value without % indicates the average rewards.
-The seeds for the Minecraft environment are fixed for the corresponding task to make a fair comparison between different models.
-Details of the tasks are provided in Section B.1.
-Policies for Atomic Skills
-We employ both language-conditioned and video-conditioned policies to learn atomic skills from the segmented demonstration videos as shown in Stage III.
-The video-conditioned policy is built upon GROOT (Cai et al., 2023b), which utilizes self-supervised learning to model instructions and behaviors.
-In the original GROOT, a fixed-length sequence of 128 frames serves as the instruction, which is encoded into a goal embedding using a conditioned variational autoencoder (C-VAE) (Kingma and Welling, 2014; Sohn et al., 2015).
-The decoder then predicts actions based on the instruction and environmental observations in an auto-regressive manner.
-The language-conditioned policy is derived from STEVE-1 (Lifshitz et al., 2023), an instruction-tuned VPT model capable of following open-ended texts or 16-frame visual instructions.
-This model is trained by first adapting the pretrained VPT model to follow commands in MineCLIP’s (Fan et al., 2022) latent space, and then training a C-VAE model to predict latent codes from the text.
-The original GROOT and STEVE-1 policies divide the training trajectories into segments of 128 frames and uniformly sample between 15 and 200 frames, respectively.
-We re-train these models using our SBD method on the segmented trajectories and compare the results to the original models to demonstrate the effectiveness of our approach.
-We retain the minimum and maximum length settings.
-The algorithm for pruning the length of each trajectory is detailed in B.4.
-Most of the original hyperparameter settings are retained as specified in the original papers.
-A complete list of modified hyperparameters is provided in Section B.2.
+See more details in Sec.˜D.1.
 Method Wood Food Stone Iron Average
@@
 $\Delta$ +1.1% +25.0% +9.8% +9.4% +11.3%
-(a) OmniJARVIS (behavior cloning)
+(a) OmniJARVIS (behavioral cloning)
 Method Wood Oak Birch Stone Iron Diamond Armor Food Average
@@
 (b) JARVIS-1 (in-context learning)
-Table 3: Success rate of two agents with their corresponding controllers trained on dataset segmented by our SBD method and the original sequential sampling (SS) method on groups of long programmatic tasks.
-All tasks are tested 30 times.
-In each group, the agent is required to obtain a certain type of items from scratch or given an iron pickaxe.
+Table 3: Success rate of two agents on long programmatic tasks.
+The goal-conditioned policies of the agents are trained on the dataset segmented by SBD and the original random splitting method.
+In each group, the agent is required to obtain a certain type of item from scratch or be given an iron pickaxe.
 For example, the diamond group includes diamond pickaxe, diamond sword, jukebox, etc.
-Task [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Average
-- 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 64.2
-Info 33.0% 52.0% 100.0% 88.0% 97.0% 32.0% 23.7 65.0% 92.0% 47.0% 15.5 21.6 87.4
-Loss 44.0% 54.0% 100.0% 94.0% 72.0% 46.0% 25.8 63.0% 95.0% 48.0% 17.9 22.7 91.8
-Both 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 93.3
-Table 4: Ablation on the components within SBD. We report the evaluation results on Minecraft atomic skills from the sequential sampling (- in the table) and SBD with different components.
-Long-term Instruction Following
-We use the hierarchical agents to model the long trajectories, which are widely used by (Driess et al., 2023; Wang et al., 2025, 2023b).
-We utilize in-context learning, as demonstrated by JARVIS-1 (Wang et al., 2024b), and imitation learning, as shown by OmniJARVIS (Wang et al., 2025), to evaluate the long-horizon instruction-following capabilities of hierarchical agents.
+3.3 Implemetation Details
+We take the Minecraft environment [18, 49] as the evaluation simulator.
+It offers a highly complex (> $10^{10^{20}}$ states), diverse, and dynamic environment ideal for testing open-ended agents requiring OOD generalization [28].
+See Appendix˜C for more information on this environment.
+We now introduce the architectures and training/inference details of the policies and agents used in Sec.˜3.1.
+Unconditional Policy and Segmentation Details
+For the unconditional policy $\pi_{\text{unconditional}}$ in Stage I, we use the pre-trained vpt-3x models [3].
+This is a foundational Transformer-XL [11] based model pretrained on large-scale YouTube data and finetuned on the early-game dataset using behavioral cloning.
+The hyperparameter GAP in Stage II is set to 18, considering the average trajectory length and semantics.
+As for the external information, we use event-based information such as mine_block:oak_log or use_item:torch available in the contractor dataset [3].
+To identify potential skill transitions, only the final event in a series of identical consecutive events is marked as positive, as it signifies a possible skill transition.
+For instance, if the agent chops a tree repeatedly and then crafts a table, the video should be split at the moment the agent mines the last block of wood.
+Further details are provided in Sec.˜D.3.
+Policies for Atomic Skills
+We employ both video-conditioned and language-conditioned policies to learn atomic skills from the segmented demonstration videos, as shown in Stage III.
+The video-conditioned policy is built upon GROOT [6], which utilizes self-supervised learning to model instructions and behaviors.
+In the original GROOT, a fixed-length sequence of 128 frames serves as the instruction, which is encoded into a goal embedding using a conditioned variational autoencoder (C-VAE) [23, 33].
+The decoder then predicts actions based on the instruction and environmental observations in an auto-regressive manner.
+The language-conditioned policy is derived from STEVE-1 [27], an instruction-tuned VPT model capable of following open-ended texts or 16-frame visual instructions.
+This model is trained by first adapting the pretrained VPT model to follow commands in MineCLIP’s [15] latent space, and then training a C-VAE model to predict latent codes from the text.
+The original GROOT and STEVE-1 policies divide the training trajectories into segments of 128 frames and uniformly sample between 15 and 200 frames, respectively.
+We re-train these models using our SBD method on the segmented trajectories and compare the results to the original models to demonstrate the effectiveness of our approach.
+We retain the minimum and maximum length settings.
+The algorithm for pruning the length of each trajectory is detailed in D.4.
+Most of the original hyperparameter settings are retained as specified in the original papers.
+A complete list of modified hyperparameters is provided in Sec.˜D.2.
+Long-term Instruction Following Agents
+We use hierarchical agents to model long trajectories, which are widely used by [12, 40, 38].
+We utilize in-context learning, as demonstrated by JARVIS-1 [39], and behavioral cloning, as shown by OmniJARVIS [40], to evaluate the long-horizon instruction-following capabilities of hierarchical agents.
 JARVIS-1 is a hierarchical agent that uses the text-conditioned STEVE-1 as its policy and pretrained vision-language models as planners.
@@
 4 Experiments and Analysis
+Task
+use
+furnace
+hunt
+sheep
+sleep
+in bed
+use
+torch
+use
+boat
+use
+bow
+collect
+stone
+collect
+seagrass
+collect
+wood
+find and
+collect wood
+collect
+dirt
+collect
+grass
+Average
+- 21.0% 26.0% 100.0% 97.0% 71.0% 30.0% 21.7 34.0% 76.0% 19.0% 14.5 9.5 63.3
+good dist. 38.0% 65.0% 90.0% 89.0% 80.0% 10.0% 24.8 43.0% 96.0% 41.0% 15.5 19.0 79.9
+Info 33.0% 52.0% 100.0% 88.0% 97.0% 32.0% 23.7 65.0% 92.0% 47.0% 15.5 21.6 85.8
+Loss (GAP=17.5) 44.0% 49.0% 100.0% 89.0% 91.0% 32.0% 27.3 65.0% 96.0% 38.0% 17.8 22.7 88.4
+Loss (GAP=18) 44.0% 54.0% 100.0% 94.0% 72.0% 46.0% 25.8 63.0% 95.0% 48.0% 17.9 22.7 90.2
+Loss (GAP=18.5) 41.0% 51.0% 100.0% 91.0% 73.0% 44.0% 26.8 57.0% 91.0% 45.0% 18.1 25.3 88.5
+Both (GAP=18) 30.0% 54.0% 100.0% 88.0% 93.0% 80.0% 26.8 51.0% 90.0% 44.0% 19.7 25.4 91.7
+Table 4: Ablation studies.
+We report the evaluation results of GROOT trained on datasets segmented by random splitting with fixed length (-), random splitting following the length distribution of SBD (good dist.), and SBD with different components and GAP values.
+A value with % indicates the average success rate, while a value without % indicates the average rewards.
+To enable a fair comparison between success rate and rewards, the average score is computed by scaling the best score to 100 and all other scores proportionally for each task.
 In our experiments, we answer the following questions:
-$\bullet$ The short segments obtained through SBD method, are they more consistent at the semantic level, and can a better atomic skill policy be trained on these segments?
-$\bullet$ Does using short videos segmented with new methods as skills perform better on long-horizon tasks?
-$\bullet$ Which component of SBD is the most important?
+$\bullet$ Are the short segments obtained through SBD more consistent semantically, and can a better goal-conditioned policy be trained on these segments?
+$\bullet$ Do improvements in goal-conditioned policies enhance the ability of hierarchical agents?
+$\bullet$ Which component of SBD is more important?
 4.1 Experimental Setups
-Environmental Setups and Datasets
-We take the Minecraft environment (Guss et al., 2019; Zheng et al., 2023) as the evaluation simulator.
-Within the Minecraft environment, we use OpenAI’s contractor dataset 7.x (early game) (Baker et al., 2022) as trajectories.
-The dataset contains offline trajectories with 68M frames with a duration of approximately 1000 hours, with at least half of the data from the first 30 minutes of the game.
-Our method generates a segmented dataset of 130k sub-trajectories using bc-early-game-3x (Baker et al., 2022) as the pretrained unconditional model.
+Training Dataset
+We use OpenAI’s contractor dataset 7.x (early game) [3] as our training dataset.
+It contains Minecraft offline trajectories with 68M frames with a duration of approximately 1000 hours, with at least half of the data from the first 30 minutes of the game.
+Our method generates a segmented dataset of 130k sub-trajectories using bc-early-game-3x [3] as the pretrained unconditional model.
 Evaluation Benchmarks
-For the atomic conditioned policies, we selected basic skills such as chop down trees and advanced skills like smelt items with furnace in Minecraft as evaluation benchmarks.
-We tested 12 different skill sets designed in MCU (Lin et al., 2023).
-Each task was tested over 100 times, except for sleep in bed and use bow, which were evaluated 10 times using the human Elo rating.
-The evaluation metrics are computed with the success rates and rewards.
-The hierarchical agents are evaluated on long-horizon programmatic tasks requiring the agents to start from an empty inventory in a new world until obtaining the final required items, such as obtain an iron pickaxe from scratch, which is usually a chain of atomic tasks.
-We split the long-horizon tasks into different groups, including wooden items, food, stones items, iron etc.
-For each group, we choose the tasks from JARVIS-1 benchmarks (Wang et al., 2024b) and evaluate each task over 10 times.
+For the goal-conditioned policies, we select basic skills such as collect wood and advanced skills like use furnace in Minecraft as the evaluation benchmark.
+We test 12 different atomic skills designed based on MCU [28].
+All tasks are tested over 100 times, except for sleep in bed and use bow, which are evaluated 10 times using human rating because they cannot be automatically verified with in-game information.
+The evaluation metrics are success rates or rewards (for finer-grained assessment of easier tasks).
+See more details in Sec.˜D.1.
+The hierarchical agents are evaluated on long-horizon programmatic tasks that require the agents to start from an empty inventory in a new world until obtaining the final required items, such as obtain an iron pickaxe from scratch, which is usually a chain of atomic tasks.
+We split the long-horizon tasks into different groups, including wooden items, food, stone items, iron etc.
+For each group, we choose the tasks from JARVIS-1 benchmarks [39] and evaluate each task over 30 times.
 4.2 Main Results
-In this section, we present the results of our method, SBD, applied to the two policies and two agents mentioned earlier.
-SBD achieves an average performance increase of 63.7% for GROOT and 52.1% for STEVE-1 compared to the original methods.
-Additionally, it enhances the performance of OmniJARVIS and JARVIS-1 by 11.3% and 20.8%.
+In this section, we first present a human analysis of the quality of the SBD-segmented videos.
+Then we present the results of SBD applied to the two policies and two agents in Sec.˜3.3 on the benchmarks in Sec.˜4.1.
+Human Analysis of Segmented Videos
+We ask 10 people to analyze the SBD-segmented videos, each with 50 randomly sampled clips: On average, 54% exhibit clear independent semantics aligned with human preference, 34% are considered acceptable, and only 12% could be better segmented.
+This further supports our method’s effectiveness.
 Short-horizon Atomic Tasks
-The results are shown in Table 2.
-Scores with % indicate success rates, while those without % represent rewards for the corresponding tasks.
-The controllers showed substantial improvements across most tasks, with an average performance enhancement of 63.7% for video-conditioned policy and 52.1% for language-conditioned policy.
+As shown in Tab.˜2, the policies showed substantial improvements across most tasks, with an average relative performance enhancement of 63.7% for GROOT and 52.1% for STEVE-1.
 Long-horizon Programmatic Tasks
-To further verify that the improvements in controllers enhance the ability of hierarchical agents, we evaluate OmniJARVIS and JARVIS-1 on programmatic tasks selected from the original papers (Wang et al., 2025, 2024b).
-As shown in Table 3, the agents have substantial improvements across most of the tasks, achieving an average performance enhancement of 11.3% for Omnijarvis and 20.8% for JARVIS-1.
-4.3 Ablation Study
+As shown in Tab.˜3, the agents have substantial improvements across most of the tasks, achieving an average relative performance enhancement of 11.3% for Omnijarvis and 20.8% for JARVIS-1.
+4.3 Ablation Studies
 (a)    Info only.
@@
 Furthermore, the similarity between the combined method and the loss-only method indicates that predictive loss is the primary factor in learning the segmentation pattern.
-Skill Videos
 Figure 3: Video Segment Examples.
@@
 The remaining three screenshots are manually selected to best illustrate the skill’s progression.
-More segments can be found in Appendix C.
-As introduced in Sections 3.2.1 and 3.2.2, the environment information and the skill boundary detection loss are important components for accurate segmentation of long videos.
-In this section, we explore the effectiveness of every component.
-So we compare the full SBD with sequential sampling (w/o loss and info), SBD w/ info only, and SBD w/ loss only.
-To evaluate the effectiveness of these configurations, we conducted an ablation study on the performance of four methods under GROOT.
-The results are presented in Table 4.
-Our findings are as follows: (i) Using external information alone significantly improves performance, with an average increase of 36.1% over the original agent.
-(ii) Using loss alone results in a better performance (an average increase of 5.0%) compared to using external information alone, highlighting the effectiveness of the core component of our method.
+More segments can be found in Appendix˜E.
+SBD Components
+As introduced in Secs.˜3.2.1 and 3.2.2, the environment information and the skill boundary detection loss are important components for accurate segmentation of long videos.
+In this section, we explore the effectiveness of each component.
+We compare the full SBD with random splitting (w/o loss and info), SBD w/ info only, and SBD w/ loss only.
+The results are presented in Tab.˜4.
+Our findings are as follows: (i) Using external information alone significantly improves performance, with a relative average increase of 35.5% over the original agent.
+(ii) Using loss alone results in a better performance (a relative average increase of 5.1%) compared to using external information alone, highlighting the effectiveness of the core component of our method.
 This demonstrates that our method achieves strong performance even without external information.
 (iii) Combining loss and external information yields the best overall performance.
-This combination outperforms the individual components in half of the tasks, suggesting that each component captures unique segmentation patterns that the other misses.
-We observe performance declines when employ both components in tasks smelt food and collect seagrass.
-This may be due to both components partitioning the trajectory at the same intervals but not at identical steps, leading to redundant splits and incorrect trajectory fragments.
+This suggests that each component captures unique segmentation patterns that the other misses.
+We observe performance declines when employing both components in tasks use furnace and collect seagrass.
+This may be due to both components splitting the trajectory at the same intervals but not identical steps, leading to redundant trajectory fragments.
+Hyperparameter sensitivity
+GAP is an important hyperparameter of SBD.
+To verify the stability of our method, we also conduct an ablation study on different values of GAP.
+As shown in Tab.˜4, our method is not sensitive to it (only 2.0% relative average performance decline).
+Improved Baseline
+A simple idea to improve random splitting is to follow a similar length distribution as SBD.
+We add this improved baseline to show that our method also catches features other than length distribution.
+As shown in Tab.˜4, while a better length distribution can improve baseline performance (by 26.2% relatively), our method still outperforms it (by 14.8% relatively).
 4.4 Visualizations
 Length Distribution
-Fig.
-2 shows the length distribution of sub-trajectories split by different methods with or without the two components of SBD.
+Fig.˜2 shows the length distribution of sub-trajectories split by different methods with or without the two components of SBD.
 We observe that both loss-only and info-only methods follow similar distributions, where the log length and frequency of sub-trajectories exhibit a normal distribution.
@@
 This observation highlights that the predictive loss is the key driver in learning the segmentation pattern, with the external information serving a more supplementary role.
-We sample one long unsegmented video segmented by our method for each skill in the datasets, such as sleep in bed and collect seagrass (see Fig. 3). More segments can be seen in Appendix C.
+Skill Videos
+We sample one long unsegmented video segmented by our method for each skill in the datasets, such as sleep in bed and collect seagrass (see Fig.˜3). More segments can be seen in Appendix˜E.
 5 Related Works
@@
 Prior work on learning from unsegmented demonstrations has proposed various methods to segment trajectories into sub-trajectories.
-For instance, some methods (Shankar and Gupta, 2020; Tanneberg et al., 2021; Kipf et al., 2019) utilize a variational autoencoder (Kingma and Welling, 2014) model to generate pairs of skill types and durations, $(g_{i},t_{i})_{i=1}^{k}$, from a given trajectory, thereby enabling segmentation.
-TACO (Shiarlis et al., 2018) adopts a weakly supervised framework that identifies task boundaries, $(t_{1},\ldots t_{k})$ based on task sketches, $(g_{1},\ldots g_{k})$, by solving a sequence alignment problem to achieve appropriate segmentation.
-BUDS (Zhu et al., 2022) constructs hierarchical task structures of demonstration sequences using a bottom-up strategy, deriving temporal segments through agglomerative clustering of the actions.
-FAST (Pertsch et al., 2025) proposes an approach based on byte-pair encoding (Gage, 1994) algorithm for segmentation and tokenization of robot action trajectories via time-series compression.
+For instance, some methods [31, 35, 24] utilize a variational autoencoder [23] model to generate pairs of skill types and boundaries, $(g_{i},t_{i})_{i=1}^{k}$, from a given trajectory.
+TACO [32] adopts a weakly supervised framework that identifies skill boundaries, $(t_{1},\ldots t_{k})$, based on skill sketches, $(g_{1},\ldots g_{k})$, by solving a sequence alignment problem.
+BUDS [52] constructs hierarchical task structures of demonstration sequences using a bottom-up strategy, deriving temporal segments through agglomerative clustering of the actions.
+FAST [29] proposes an approach based on the byte-pair encoding [16] algorithm for segmentation and tokenization of robot action trajectories.
 Option Learning
-Option learning (Sutton et al., 1999) is a framework designed to learn temporal abstractions (i.e., skills or primitives) directly from the environment’s reward signal, primarily through online reinforcement learning.
-Recent approaches (Bacon et al., 2017; Bagaria and Konidaris, 2019; Klissarov and Precup, 2021) rely on policy gradient methods to learn the option policies.
-This framework requires reward signals, and DIAYN (Eysenbach et al., 2019) proposes a method to generate a reward function adaptively in the absence of reward signals, resulting in the unsupervised emergence of diverse skills.
+Option learning [34] is a framework to learn temporal abstractions (i.e., options or skills) directly from the environment’s reward signal, primarily through online reinforcement learning.
+Recent approaches [1, 2, 25] rely on policy gradient methods to learn the option policies.
+DIAYN [14] proposes a method to generate a reward function adaptively in the absence of reward signals, resulting in the unsupervised emergence of diverse skills.
 Event Segmentation Theory
-In neuroscience, episodic memory is the memory of everyday events consisting of short slices of experience (Conway, 2009).
-Event Segmentation Theory (EST) (Zacks et al., 2007) offers a theoretical perspective on how the neurocognitive system splits a long flow of memory into short events.
+In neuroscience, episodic memory is the memory of everyday events consisting of short slices of experience [10].
+Event Segmentation Theory (EST) [44] offers a theoretical perspective on how the neurocognitive system splits a long flow of memory into short events.
 According to EST, observers build event models of the current situation to generate predictions of future perceptual input.
-When errors in predictions arise, an event boundary is perceived, causing the event model to be reset and rebuilt.
+When errors in predictions arise, an event boundary is perceived, and the event model is reset and rebuilt.
 Our method employs a pretrained model to predict the agent’s action based on past observations and uses the predictive loss as an indicator to detect skill boundaries.
 Agents in Minecraft
-Many agents have been developed to interact with Minecraft environments (Zhao et al., 2024a; Jiang et al., 2025).
-For short-horizon tasks, methods typically employ imitation or reinforcement learning, as seen in works like VPT (Baker et al., 2022), which annotated a large YouTube Minecraft video dataset with actions and trained the first foundation agent in the domain using behavior cloning, and its derivatives (Lifshitz et al., 2023; Cai et al., 2023b; Yuan et al., 2024; Cai et al., 2023a, 2024b).
-For long-horizon, programmatic tasks, large language models (LLMs) are used as planners combined with skill policies (Wang et al., 2023b; Zhou et al., 2024; Yuan et al., 2023; Qin et al., 2024; Wang et al., 2024b; Li et al., 2024), and some methods (Wang et al., 2024b; Zhu et al., 2023; Wang et al., 2024a) employ explicit memory mechanisms to further enhance the long-horizon capabilities of LLM agents.
-Additionally, recent advances (Wang et al., 2025; Zhao et al., 2024b) have explored using end-to-end visual-language models (VLMs) to directly follow human instructions.
-Finally, some research (Guo et al., 2024; Zhang et al., 2023) focuses on open-ended creative tasks such as building and decoration, which often cannot be directly defined by rule-based rewards.
-6 Conclusions
-In this paper, we propose a novel temporal video segmentation algorithm to address the challenges of open-world skill discovery from unsegmented demonstrations.
+Many agents have been developed to interact with Minecraft environments [47, 21].
+For short-horizon tasks, methods typically employ behavioral cloning or reinforcement learning, as seen in works like VPT [3], which annotates a large Minecraft YouTube video dataset with actions and trains the first foundational agent in the domain, and its derivatives [27, 6, 43, 5, li2025jarvis].
+For long-horizon, programmatic tasks, large language models (LLMs) are used as planners combined with skill policies [38, 50, 42, 30, 39, 26, ui-tars-15-seed], and some methods [39, 51, 36] employ explicit memory mechanisms to enhance the long-horizon capabilities of LLM agents.
+Additionally, recent advances [40, 48] have explored using end-to-end vision-language models (VLMs) to directly follow human instructions.
+Finally, some research [17, 45] focuses on open-ended creative tasks such as building and decoration.
+6 Conclusion
+In this paper, we propose a novel temporal video segmentation algorithm to address the challenges of open-world skill discovery from unsegmented demonstration videos.
 To generate a segmented dataset of video clips with independent skills, our algorithm detects the potential skill boundaries based on the predictive loss of a pretrained action-prediction model.
-External boundary indicators can also be incorporated in our method.
-The method does not require any manual annotations and can be employed directly on massive internet gameplay videos.
-7 Limitations and Future work
-The current algorithm is relatively time-consuming, requiring approximately 1 minute of processing on an NVIDIA RTX 4090Ti for every 5-minute video, due to the need to predict actions and compute loss across the entire trajectory.
-As a result, we have not tested our method on larger datasets.
-Moreover, evaluating the method requires reprocessing the entire dataset and retraining the controllers, so we have not conducted an extensive ablation study on pretrained model selection or hyperparameter tuning.
-Additionally, we observe that the algorithm becomes unstable in action-intensive scenarios (e.g., combat situations), often generating many short segments.
-Future work could focus on improving the efficiency and stability of the algorithm, as well as exploring more effective evaluation methods.
-We also anticipate the potential application of our method to larger datasets and other open-world video games beyond Minecraft.
+It does not require any manual annotations and can be employed directly on massive internet gameplay videos.
 Acknowledgement
 This work is funded in part by the National Science and Technology Major Project 2022ZD0114902. We thank a grant from CCF-Baidu Open Fund.
-Appendix A Proofs
-In this section, we provide a detailed proof of Theorem 3.4, regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
+Supplementary Material
+Appendix A Discussion and Future work
+Computational Cost
+Our method takes approximately 1 minute to process a 5-minute video on an NVIDIA RTX 3090Ti, due to the need to predict actions across the entire trajectory.
+Since data preprocessing is a one-time cost, spending only 1/20 of the video duration with 4 GPUs is acceptable.
+Also, it is more efficient than those methods relying on human or VLM annotations while ensuring quality.
+Failure Case Due to Occasional Assumption Violation
+Our method is occasionally unstable in action-intensive scenes like combat, where performance drops on task combat spiders (29% to 20%).
+See Appendix˜E for visualization of this failure case.
+This is likely due to a violation of Assumption˜3.2 (Skill Confidence), as indicated by sharp prediction loss fluctuations.
+However, such issues are not observed elsewhere.
+Since overly short segments (< 5 frames) comprise less than 10% of all videos, overall effectiveness remains unaffected.
+Choice of Loss Term
+Currently our method detects changes in $P(a_{t}|o_{1:t})$.
+A possible alternative is to use $P(o_{t+1}|o_{1:t})$.
+We don’t use it because it does not directly reflect skill changes as $P(a_{t}|o_{1:t})$ does, and current models [50] for predicting future observations suffer from hallucinations.
+We leave the attempt of this alternative to future work.
+Scalability and Broader Applicability
+Since our method is label-free, it has the potential to segment and train on the numerous gameplay trajectories on YouTube.
+We also anticipate the potential application of our method to other larger datasets and other open-world video games beyond Minecraft.
+Appendix B Proofs
+In this section, we provide a detailed proof of Theorem˜3.4, regarding the bounds of relative predictive probability under scenarios of skill transition and non-transition.
 We prove the lower bound and upper bound respectively in the following two subsections.
-A.1 Proof of Upper Bound Under Skill Non-Transition
+B.1 Proof of Upper Bound Under Skill Non-Transition
 $P(\frac{P(a_{t+1}|o_{1:t+1})}{(\prod_{i=1}^{t}P(a_{i}|o_{1:i}))^{1/t}}>\frac{(K-1)c}{K})$ $>P(P(a_{t+1}|o_{1:t+1})>\frac{(K-1)c}{K})$
 […]
-$>P(\pi_{t}(a_{t+1}|o_{1:t+1})>c)$ (Assumption 3.1)
+$>P(\pi_{t}(a_{t+1}|o_{1:t+1})>c)$ (Assumption˜3.1)
 $=P(\pi_{t+1}(a_{t+1}|o_{1:t+1})>c)$
-$>1-\delta$ (Assumption 3.2)
-A.2 Proof of Lower Bound Under Skill Transition
+$>1-\delta$ (Assumption˜3.2)
+B.2 Proof of Lower Bound Under Skill Transition
 We first magnify the likelihood ratio between the next action and the average action history,
@@
 $<\frac{P(\pi_{t+1}=\pi_{t}|o_{1:t+1})\pi_{t}(a_{t+1}|o_{1:t+1})+P(\pi_{t+1}\neq\pi|o_{1:i})}{\frac{K-1}{K}(\prod_{i=1}^{t}\pi_{i-1}(a_{i}|o_{1:i}))^{1/t}}$ (LABEL:eq:transition and
-Assumption 3.1)
-$<\frac{\pi_{t}(a_{t+1}|o_{1:t+1})+\frac{1}{K}}{\frac{K-1}{K}(\prod_{i=1}^{t}\pi_{i-1}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption 3.1)
+Assumption˜3.1)
+$<\frac{\pi_{t}(a_{t+1}|o_{1:t+1})+\frac{1}{K}}{\frac{K-1}{K}(\prod_{i=1}^{t}\pi_{i-1}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption˜3.1)
 $=\frac{K}{K-1}\frac{\pi_{t}(a_{t+1}|o_{1:t+1})}{\prod_{i=1}^{t}\pi_{t}(a_{i}|o_{1:i}))^{1/t}}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}$
-$<\frac{Km}{2(K-1)}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption 3.3)
+$<\frac{Km}{2(K-1)}+\frac{1}{(K-1)\prod_{i=1}^{t}\pi_{i}(a_{i}|o_{1:i}))^{1/t}}$ (Assumption˜3.3)
 Therefore,
@@
 $>\prod_{i=1}^{t}P(\pi_{i}(a_{i}|o_{1:i})>c)$
-$>(1-\delta)^{t}>1-t\delta$ (Assumption 3.2)
-Appendix B Experiment Details
-In this section, we provide our experiment details including benchmark, training details for GROOT and STEVE-1, how we use event-based information as external auxiliary information, and the length pruning algorithm for sub-trajectories.
-B.1 Tasks in the Early Game Benchmark
-Minecraft SkillForge is a diverse benchmark of 30 short atomic tasks that can comprehensively evaluate the mastery of atomic skills by agents in Minecraft.
-We choose 10 early game tasks from the original benchmark which the original agent can already handle at a basic level because our method is an improvement on dataset processing instead of a new model architecture, so it cannot help agents learn new skills that they are completely incapable of acquiring previously.
-Besides, we add the task "use torch", which is also a useful skill in Minecraft; and we add an enhanced version of the task "collect wood", which requires the agent to start from the plains instead of the forest, aiming to test its ability to explore and find trees.
-Details of the 12 tasks in our early game benchmark are shown in Table 5.
-For each task, we include the evaluation metric and a brief description of what the task is.
-For tasks "Sleep" and "Use bow", GROOT performs very well on the original metric, so we design a manual metric that better assesses its actual performance.
+$>(1-\delta)^{t}>1-t\delta$ (Assumption˜3.2)
+Appendix C Minecraft Environment
+Minecraft is an extremely popular sandbox game that allows players to freely create and explore their world.
+This game has infinite freedom, allowing players to change the world and ecosystems through building, mining, planting, combating, and other methods.
+It is precisely because of this freedom that Minecraft becomes an excellent AI testing benchmark.
+In this game, AI agents need to face situations that are highly similar to the real world, making judgments and decisions to deal with various environments and problems.
+By using Minecraft, AI researchers can more conveniently simulate various complex and diverse environments and tasks, thereby improving the practical value and application of AI technology.
+Our Minecraft environment is a hybrid between MineRL [18] and the MCP-Reborn (https://github.com/Hexeption/MCP-Reborn) Minecraft modding package.
+Unlike the regular Minecraft game, in which the server (or the "world") always runs at 20Hz while the client’s rendering speed can typically reach 60-100Hz, the frame rate is fixed at 20 fps for the client in our experiments.
+The action and observation spaces in our environment are identical to what a human player can operate and observe on their device when playing the game.
+These details will be further explained in subsequent subsections.
+C.1 Minecraft Game World Setting
+We choose Minecraft version 1.16.5’s survival mode as our experiment platform.
+In this mode, the agent may encounter situations that result in its death, such as being burned by lava or a campfire, getting killed by hostile mobs, or falling from great heights.
+When this happens, the agent will lose all its items and respawn at a random location near its initial spawn point within the same Minecraft world or at the last spot it attempted to sleep.
+Importantly, even after dying, the agent retains knowledge of its previous deaths and can adjust its actions accordingly since there is no masking of policy state upon respawn.
+C.2 Observation Space
+The environmental observation space consists of two parts.
+The first part is the raw pixels from the Minecraft game that players would see, including overlays such as the hotbar, health indicators, and animations of a moving hand in response to attack or "use" actions.
+The rendering resolution of Minecraft is 640x360; however, in our experiments, we resize images to 128x128 for better computational efficiency while maintaining discernibility.
+The second part includes auxiliary information about the agent’s current environment, such as its location and weather conditions.
+Human players can obtain this information by pressing F3.
+The specific observation details we include are shown in Tab.˜5.
+Sources Shape Description
+pov (raw pixel)
+(640, 360, 3)
+to (128,128,3)
+Ego-centric RGB frames.
+player_pos (5,) The coordinates of (x,y,z), pitch, and yaw of the agent.
+location_stats (9,)
+The environmental information of the agent’s current position,
+including biome_id, sea_level, can_see_sky, is_raining etc.
+inventory (36,)
+The items in the current inventory of the agent, including
+the type and corresponding quantity of each item in each slot.
+If there is no item, it will be displayed as air.
+equipped_items (6,)
+The current equipment of the agent, including mainhand, offhand,
+chest, feet, head, and legs slots. Each slot contains type, damage,
+and max_damage information.
+event_info (5,)
+The events that occur in the current step of the game, including
+pick_up (picking up items), break_item (breaking items),
+craft_item (crafting items using a crafting table or crafting grid),
+mine_block (mining blocks by suitable tools), and
+kill_entity (killing game mobs).
+Table 5: The observation space we use in Minecraft.
+During the actual inference process, the controllers (VPT [3], GROOT [6], STEVE-1 [27]) only perceive the raw pixels.
+The agents (JARVIS-1 [39], OmniJarvis [40]) can access auxiliary information from the environment to generate the text condition of the controller.
+C.3 Action Space
+Our action space includes almost all actions directly available to human players, such as keypresses, mouse movements, and clicks.
+It consists of two parts: the mouse and the keyboard.
+When in-game GUIs are not open, the mouse movement is responsible for changing the player’s camera perspective.
+When a GUI is open, it moves the cursor.
+The left and right buttons are responsible for attacking and using items.
+The keyboard is mainly responsible for controlling the agent’s movement.
+We use the same joint hierarchical action space as VPT [3], which combines button space and camera space.
+Button space encodes all combinations of possible keyboard operations (excluding mutually exclusive actions such as forward and back) and a flag indicating whether the mouse is used, resulting in a total of 8461 candidate actions.
+The camera space discretizes the range of one mouse movement into 121 actions.
+Therefore, the action head of the agent is a multi-classification network with 8461 dimensions and a multi-classification network with 121 dimensions.
+We also filter out frames (both the observation and action) with null action as VPT [3].
+In addition, we abstract the crafting and smelting actions with GUI into functional binary actions, which are the same as MineDojo (Fan et al., 2022).
+These advanced actions are only used by the agents (JARVIS-1 [39], OmniJarvis [40]).
+The detailed action space is described in Tab.˜6.
+Index Action Human Action Description
+1 Forward key W Move forward.
+2 Back key S Move backward.
+3 Left key A Strafe left.
+4 Right key D Strafe right.
+5 Jump key Space Jump. When swimming, keeps the player afloat.
+6 Sneak key left Shift Slowly move in the current direction of movement.
+7 Sprint key left Ctrl Move quickly in the direction of the current motion.
+8 Attack left Button Destroy blocks (hold down); Attack entity (click once).
+9 Use right Button Interact with the block that the player is currently looking at.
+10 hotbar.[1-9] keys 1 - 9 Selects the appropriate hotbar item.
+11 Yaw move Mouse X Turning; aiming; camera movement.Ranging from -180 to +180.
+12 Pitch move Mouse Y Turning; aiming; camera movement.Ranging from -180 to +180.
+13 Equip - Equip the item in the main hand from the inventory.
+14 Craft - Execute a crafting recipe to obtain a new item.
+15 Smelt - Execute a smelting recipe to obtain a new item.
+Table 6: The action space we use in Minecraft.
+Appendix D Experiment Details
+In this section, we provide our experiment details, including benchmark, training details for GROOT and STEVE-1, how we use event-based information as external auxiliary information, and the length pruning algorithm for sub-trajectories.
+D.1 Minecraft Skill Benchmark
+MCU [28] is a diverse benchmark that can comprehensively evaluate the mastery of atomic skills by agents in Minecraft.
+Since our method is an improvement on dataset processing instead of a new model architecture, it cannot help agents learn new skills that they are completely incapable of acquiring previously.
+Therefore, we choose 10 early game skills from the original benchmark that the original agent can already grasp at a basic level.
+Besides, we add use torch, which is also a useful skill in Minecraft.
+We also add find and collect wood, an enhanced version of the skill collect wood, which requires the agent to start from the plains instead of the forest, aiming to test its ability to explore and find trees.
+Details of the 12 skills in our early game benchmark are shown in Tab.˜7.
+For each skill, we include the evaluation metric and a brief description of what the skill is.
+For skills "Sleep" and "Use bow", GROOT performs very well on the original metric, so we design a manual metric that better assesses its actual performance.
 STEVE-1 can take text as prompts, which are also listed in the table.
-For some of the tasks, it is tricky to find proper text prompts, so we use visual prompts instead.
-Task Metric Description Text Prompt for STEVE-1
-[Uncaptioned image] Craft item cooked mutton Given a furnace and some mutton and coal, craft a cooked mutton. -
-[Uncaptioned image] Kill entity sheep Summon some sheep before the agent, hunt the sheep. -
-[Uncaptioned image] STEVE-1: Use item white bed. GROOT: Sleep in bed properly. Given a white bed, sleep on it. Sleep in bed.
-[Uncaptioned image] Use item torch Give some torches, use them to light up an area. Time is set at night. Use a torch to light up an area.
-[Uncaptioned image] Use item birch boat Given a birch boat, use it to travel on water. Biome is ocean. -
-[Uncaptioned image] STEVE-1: Use item bow. GROOT: Use item bow 20%, Shoot in distance 40%, take aim 40% Given a bow and some arrows, shoot the sheep summoned before the agent. -
-[Uncaptioned image] Mine block stone (cobblestone, iron, coal, diamond) Given an iron pickaxe, collect stone starting from cave. Night vision is enabled. Collect stone.
-[Uncaptioned image] Mine block seagrass (tall seagrass, kelp) Given an iron pickaxe, collect seagrass starting from ocean. -
-[Uncaptioned image] Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, collect wood starting from forest. Chop a tree.
-[Uncaptioned image] Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, find and collect wood starting from plains. Chop a tree.
-[Uncaptioned image] Mine block dirt (grass block) Given an iron pickaxe, collect dirt starting from plains. Collect dirt.
-[Uncaptioned image] Mine block grass (tall grass) Given an iron pickaxe, collect grass starting from plains. Collect grass.
-Table 5: Details of 12 atomic tasks in our early game benchmark for testing GROOT and STEVE-1.
-B.2 Training Details
-The modified hyperparameters for GROOT and STEVE-1 are listed in Table 6.
+For some of the skills, it is tricky to find proper text prompts, so we use visual prompts instead.
+Skill Metric Description Text Prompt for STEVE-1
+Use furnace Craft item cooked mutton Given a furnace and some mutton and coal, craft a cooked mutton. -
+Hunt Sheep Kill entity sheep Summon some sheep before the agent, hunt the sheep. -
+Sleep in bed STEVE-1: Use item white bed. GROOT: Sleep in bed properly. Given a white bed, sleep on it. Sleep in bed.
+Use torch Use item torch Give some torches, use them to light up an area. Time is set at night. Use a torch to light up an area.
+Use boat Use item birch boat Given a birch boat, use it to travel on water. The biome is ocean. -
+Use bow STEVE-1: Use item bow. GROOT: Use item bow 20%, Shoot in distance 40%, take aim 40% Given a bow and some arrows, shoot the sheep summoned before the agent. -
+Collect stone Mine block stone (cobblestone, iron, coal, diamond) Given an iron pickaxe, collect stone starting from cave. Night vision is enabled. Collect stone.
+Collect seagrass Mine block seagrass (tall seagrass, kelp) Given an iron pickaxe, collect seagrass starting from ocean. -
+Collect wood Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, collect wood starting from forest. Chop a tree.
+Find and collect wood Mine block oak (spruce, birch, jungle, acacia) log Given an iron pickaxe, find and collect wood starting from plains. Chop a tree.
+Collect dirt Mine block dirt (grass block) Given an iron pickaxe, collect dirt starting from plains. Collect dirt.
+Collect grass Mine block grass (tall grass) Given an iron pickaxe, collect grass starting from plains. Collect grass.
+Table 7: Details of 12 atomic skills in our Minecraft skill benchmark for testing GROOT and STEVE-1.
+D.2 Training Details
+The modified hyperparameters for GROOT and STEVE-1 are listed in Tab.˜8.
 We adjust parallel GPUs, gradient accumulation batches, and batch sizes to better align with our available computing resources.
@@
 All models are trained parallelly on four NVIDIA RTX 4090Ti GPUs.
-We follow the same policy training pipeline as the original paper except for the modified hyperparameters mentioned above.
-GROOT is trained under three epochs of our dataset.
-STEVE-1, however, is only trained under 1.5 epochs of our dataset, because we observe that the model starts to overfit on the dataset if it is trained further.
+We follow the same policy training pipeline as the original paper, except for the modified hyperparameters mentioned above.
+GROOT is trained for three epochs of our dataset.
+STEVE-1, however, is only trained for 1.5 epochs of our dataset, because we observe that the model starts to overfit on the dataset if it is trained further.
 Hyperparameter Value
@@
 Table 8: Modified hyperparameters for training controllers. (Left) GROOT. (Right) STEVE-1.
-B.3 Event-Based Information
+D.3 Event-Based Information
 Within the Minecraft environment, we focus on events in the categories "use item", "mine block", "craft item", and "kill entity", as they reflect the player’s primary activities.
 Multiple events may occur simultaneously and are recorded as a set.
-Only the final step of a repeating sequence of event sets is marked as positive.
+Only the final step of a repeating sequence of sets of events is marked as positive.
 For example, in the sequence: ("use item iron pickaxe", "mine block iron ore"), ("use item iron pickaxe", "mine block iron ore"), ("use item iron pickaxe", "mine block diamond ore"), ("use item torch"), ("use item torch"), the second, third, fifth steps will be marked as positive.
-Additionally, for any step $t$ that includes a "kill entity" event, we mark $t+16$ as positive instead of $t$, because there will be a short death animation that follows the entity’s death and we want it to be included in the same segment.
-B.4 Length Pruning Algorithm
-STEVE-1 forces a minimum and maximum length of trajectories in its method so we need a length pruning algorithm.
+Additionally, for any step $t$ that includes a "kill entity" event, we mark $t+16$ as positive instead of $t$, because there will be a short death animation that follows the entity’s death, and we want it to be included in the same segment.
+D.4 Length Pruning Algorithm
+STEVE-1 forces a minimum and maximum length of trajectories in its method, so we need a length pruning algorithm.
 In our implementation, the length of each trajectory is pruned as follows: If a trajectory is too short, it is merged with subsequent trajectories until it meets the minimum length requirement.
@@
 There might be more sophisticated strategies, but we use this straightforward algorithm for simplicity.
-Appendix C Examples of Skill Videos
-We sample one video segmented by our method for each skill in the benchmark, presenting each video with five screenshots.
+Appendix E Examples of Skill Videos
+We sample one video segmented by our method for each skill in the benchmark and an extra video showing a failure case, presenting each video with five screenshots.
 The first and last screenshots correspond to the first and last frames of the video, while the other three are manually selected to best illustrate the progression of the skill.
@@
   • collect grass
+  • combat spider (failure case: overly short)